### v1

# Affine Chord Sobolev Inequalities from Integral Geometry

Abstract.
Functional analogues of isoperimetric inequalities for chord power integrals are established for $\alpha>0$, together with a complete characterization of the equality cases. For $\alpha\in(0,n)$, these inequalities—referred to as chord Sobolev inequalities— serves as a counterpart to the fractional Sobolev inequality of Almgren and Lieb, which corresponds to the functional isoperimetric inequality for chord power integrals with $\alpha\in(-1,0)$.
Moreover, affine isoperimetric inequalities for chord power integrals are derived, that are significantly stronger than their Euclidean counterparts. Extending the framework to functional and affine settings, functional extensions of chord power integrals and radial mean bodies are introduced. The inclusion relations for classical radial mean bodies due to Gardner and Zhang are extended to both log-concave and $s$-concave function, and the extensions are shown to be sharp.
1. Introduction
For $\alpha>-1$, the chord power integral of a convex body $K\subset\mathbb{R}^{n}$ (a convex compact set with non-empty interior), is defined by
$I_{\alpha+1}(K)=\int_{{\rm Aff}(n,1)}|K\cap l|^{\alpha+1}dl,$
where ${\rm Aff}(n,1)$ is the affine Grassmannian consistent of $1$-dimensional lines in $\mathbb{R}^{n}$, and $dl$ is the Haar measure on it. The quantity $|K\cap l|$ represents the length of the intersection. The chord power integral has attracted increasing attention in recent years, mainly in connection with Minkowski-type problems (see for example [10, 17, 40, 48, 49]). In this paper, we study inequalities from the viewpoint of chord power integrals, while previous studies on related inequalities approached their problems from different perspectives.
The isoperimetric inequalities for chord power integrals (cf. [46]) state that, given a convex body $K\subset\mathbb{R}^{n}$,
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\geq\Big(\frac{|K|}{|B^{n}|}\Big)^{\frac{n+\alpha}{n}},$ (1.1)
for $\alpha\in(-1,0)\cup(n,\infty)$, and
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\leq\Big(\frac{|K|}{|B^{n}|}\Big)^{\frac{n+\alpha}{n}},$ (1.2)
for $\alpha\in(0,n)$. Equality holds in both cases if and only if $K$ is a ball. Throughout, $|K|$ denotes the volume of $K$, that is, the $n$-dimensional Lebesuge measure of $K$, and $B^{n}$ stands for the Euclidean unit ball in $\mathbb{R}^{n}$ with volume $\omega_{n}$.
For $\alpha\in(-1,0)$, the chord power integral $I_{\alpha+1}(K)$ is also referred to as the fractional perimeter, which is closely related to the fractional Sobolev norm (cf. [13, 32, 33]):
$\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,$
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a measurable function, and $|\cdot|$ denotes the Euclidean norm in $\mathbb{R}^{n}$. Replacing $f$ with the characteristic function $\operatorname{1}_{K}$ of a convex body $K\subset\mathbb{R}^{n}$, which takes value $1$ on $K$ and 0 elsewhere, yields
$I_{\alpha+1}(K)=\frac{-\alpha(\alpha+1)}{n\omega_{n}}\int_{K}\int_{K^{c}}\frac{1}{|x-y|^{n-\alpha}}dxdy,$ (1.3)
which follows from the Blaschke-Petkantschin formula (cf. [46, Theorem 7.2.7]).
Almgren and Lieb [1] established the Pólya–Szegö inequalities for fractional Sobolev norms, and the equality case was subsequently characterized by Frank and Seiringer [13]. A direct consequence is the fractional Sobolev inequality: for $\alpha\in(-1,0)$ and $f\in W^{-\alpha,1}(\mathbb{R}^{n})$,
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,$ (1.4)
where $\|f\|_{p}$ denotes the $L^{p}$ norm of $f$ and $W^{-\alpha,1}(\mathbb{R}^{n})$ denotes the fractional Sobolev space of $L^{1}$ functions with finite $\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}$. Equality holds in (1.4) if and only if $f$ is a constant multiple of the characteristic function of a ball. The best constant $\sigma_{n,\alpha}$ is given by
$\sigma_{n,\alpha}=\frac{nI_{\alpha+1}(B^{n})}{|\alpha|(\alpha+1)\omega_{n}^{\frac{\alpha}{n}}}=\frac{n2^{\alpha}\pi^{\frac{n-\alpha-1}{2}}\Gamma(\frac{\alpha+1}{2})\Gamma(\frac{n}{2}+1)^{\frac{\alpha}{n}}}{|\alpha|\Gamma(\frac{n+\alpha}{2}+1)},$ (1.5)
where $\Gamma$ denotes the gamma function. The constant can be obtained from the results in [12] as described in [16].
Given the sharp fractional Sobolev inequality introduced above—together with its best constant and equality cases—and the geometric isoperimetric inequality for chord power integrals, it becomes natural to regard (1.4) as a functional analogue of (1.1) for $\alpha\in(-1,0)$. More precisely, replacing $f$ with the characteristic function $\operatorname{1}_{K}$ of a convex body $K\subset\mathbb{R}^{n}$ reduces the inequality to the isoperimetric inequality for $I_{\alpha+1}(K)$ with $\alpha\in(-1,0)$.
Building on this correspondence, it is natural to ask whether the isoperimetric inequalities established for $\alpha\in(0,n)$ and $\alpha>n$ also admit functional counterparts.
Question. Does there exist a functional inequality for a suitable class of functions $f$ whose geometric instance, corresponding to $f=\operatorname{1}_{K}$, recovers the isoperimetric inequalities for chord power integrals with $\alpha>0$?
Noting that for $\alpha>0$, the quantity $I_{\alpha+1}(K)$ also admits a double integral representation, as shown in [40, Lemma 3.3]:
$I_{\alpha+1}(K)=\frac{\alpha(\alpha+1)}{n\omega_{n}}\int_{K}\int_{K}\frac{1}{|x-y|^{n-\alpha}}dxdy.$
This representation suggests a connection with the Hardy-Littlewood-Sobolev inequality (HLS inequality) established by Lieb [30]: for $0<\alpha<n$ and non-negative $f\in L^{p}(\mathbb{R}^{n})$ with $p=2n/(n+\alpha)$,
$\gamma_{n,\alpha}\|f\|_{\frac{2n}{n+\alpha}}^{2}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{f(x)f(y)}{|x-y|^{n-\alpha}}dxdy.$ (1.6)
Equality in (1.6) occurs if and only if $f(x)=a(1+\lambda|x-x_{0}|^{2})^{-(n+\alpha)/2}$ for $x\in\mathbb{R}^{n}$ with $a\geq 0,\lambda>0$ and $x_{0}\in\mathbb{R}^{n}$. This observation leads to the unexpected conclusion that the HLS inequality is not the functional counterpart of (1.2). On further inspection, the essential obstruction appears to be the $2$-homogeneity of the HLS inequality in $f$. Motivated by the 1-homogeneous structure arising in the case $\alpha\in(-1,0)$, we turn to a formulation that reflects the intrinsic homogeneity underlying the geometric setting. This leads naturally to the following 1-homogeneous analogue of (1.6).
Theorem 1 (Chord Sobolev inequalities).
For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy,$
where the sharp constant $\sigma_{n,\alpha}$ is given by (1.5). Equality holds if and only if $f$ is a constant multiple of the characteristic function of a ball.
With the double integral appearing in Theorem 1 and the fractional Sobolev norm (1.3), we can naturally extend the notion of the chord power integral to functions $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ by setting
$I_{\alpha+1}(f)=\frac{-\alpha(\alpha+1)}{2n\omega_{n}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,\quad\quad\alpha\in(-1,0),$
and
$I_{\alpha+1}(f)=\frac{\alpha(\alpha+1)}{n\omega_{n}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy,\quad\quad\alpha>n.$
A more unified formulation can be written as
$I_{\alpha+1}(f)=\int_{0}^{\infty}I_{\alpha+1}(\{f\geq r\})dr,$
which follows from the coarea formula when $\alpha\in(-1,0)$ and from the layer-cake representation of $\min\{f(x),f(y)\}$ when $\alpha>n$. Here $f$ is typically assumed to be an integrable log-concave function, although the log-concavity assumption is not required in general. For instance, when $\alpha\in(-1,0)$, it suffices that $f$ is a non-negative function in the corresponding fractional Sobolev space.
The inequality in Theorem 1 is a functional counterpart of the classical isoperimetric inequality (1.2). It is invariant under translations and rotations, but not under volume-preserving linear transformations. Affine inequalities, which remain invariant under translations and volume-preserving linear transformations, have proved to be very powerful, as they often strengthen the corresponding Euclidean inequalities. A famous example is the Petty projection inequality for convex bodies in $\mathbb{R}^{n}$. It strengthens the Euclidean isoperimetric inequality and also directly implies it. See, for example, [14, 45], and see [7, 19, 23, 29, 36, 37, 39, 47, 51] for extensions in the $L^{p}$ Brunn–Minkowski theory and for more general sets. Recently, Milman and Yehudayoff [42] established isoperimetric inequalities for affine quermassintegrals. These new inequalities are significantly stronger than the isoperimetric inequalities for the classic quermassintegrals, thereby settling a long-standing conjecture of Lutwak [35].
More recently, Haddad and Ludwig [25] established the following sharp affine fractional Sobolev inequality, which is significantly stronger than the classical inequality (1.4).
Affine fractional Sobolev inequalities: For $1<\alpha<0$ and $f\in W^{-\alpha,1}(\mathbb{R}^{n})$, that is, $f$ with finite $\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}$,
$\begin{split}2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha}\Big\|\frac{f(\cdot+tu)-f(\cdot)}{t}\Big\|_{1}dt\Big)^{\frac{n}{\alpha}}\Big)^{\frac{\alpha}{n}}\\ \leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy.\end{split}$ (1.7)
There is equality in the first inequality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid. There is equality in the second inequality for radially symmetric functions.
In the limit as $\alpha\rightarrow-1^{+}$, the inequality (1.7) recovers the well-known affine Sobolev inequality established by Zhang [51], which is an affine version of the classic $L^{1}$ Sobolev inequality. For a Sobolev function $f$, Zhang [51] defined the polar projection body $\Pi^{*}f$ by
$\|\xi\|_{\Pi^{*}f}=\frac{1}{2}\int_{\mathbb{R}^{n}}|\nabla f(x)\cdot\xi|dx,~~\xi\in\mathbb{S}^{n-1},$
where $\|\cdot\|_{K}$ denotes the gauge function of $K$. Wang [47] extended the definition to functions of bounded variation and obtained the generalized Sobolev-Zhang inequality. The extended version corresponds to the functional form of the Petty projection inequality. Using (1.7), Haddad and Ludwig derived the fractional Petty projection inequalities by taking $f=\operatorname{1}_{K}$.
Motivated by the same central ideas as in their earlier work [25], Haddad and Ludwig [26, 27] established the affine fractional $L^{p}$ Sobolev inequalities as well as affine HLS inequalities. For further results on affine functional inequalities, we refer the reader to [9, 11, 22].
The following Theorem 2 presents an affine version of our functional isoperimetric inequality for chord power integrals, and thus strengthens Theorem 1.
Theorem 2.
For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}\Big)^{\frac{\alpha}{n}}$ (1.8)
$\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
There is equality in the first inequality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid. There is equality in the second inequality if $f$ is radially symmetric.
In Section 4, we introduce an anisotropic version of the right side of (1.8), which gives rise to a new star-shaped set ${\rm L}_{\alpha}f$, defined via its radial function by
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,~~u\in\mathbb{S}^{n-1}.$
With the definition, the first inequality in Theorem 2 can be written as
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$ (1.9)
Since both sides of (1.9) are invariant under translations of $f$ and
${\rm L}_{\alpha}(f\circ\phi^{-1})=\phi{\rm L}_{\alpha}f$
for every volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, it follows that inequality (1.9) is indeed affine invariant.
According to (1.1) and (1.2) the isoperimetic inequality for $I_{\alpha+1}(K)$ reverses when $\alpha>n$. Theorem 1 establishes the functional analogue for $\alpha\in(0,n)$. However, the case $\alpha>n$ does not admit a straightforward reversal, which will be discussed in detail in Section 6.
The next result presents the affine inequality we obtain for $\alpha>n$, where ${\rm L}_{\alpha}f$ has been introduced above and $L^{\infty}(\mathbb{R}^{n})$ is the space of essentially bounded measurable functions, equipped with the essential supremum norm $\|\cdot\|_{\infty}$. By eliminating the intermediate term, it serves as the functional analogue of (1.1) for $\alpha>n$.
Theorem 3.
Let $\alpha>n$ and $f$ be a non-zero, non-negative function in $L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$. Then
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
There is equality in the first inequality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid. There is equality in the second inequality if $f$ is radially symmetric.
In Section 7 and Section 8, we relate the set ${\rm L}_{\alpha}f$ to radial mean bodies originally introduced by Gardner and Zhang [18]. For $\alpha>-1$ and a convex body $K\subset\mathbb{R}^{n}$, the radial $\alpha$-th mean body ${\rm R}_{\alpha}K$ is defined via its radial function by
$\rho_{{\rm R}_{\alpha}K}(u)=\Big(\frac{1}{|K|}\int_{K}\rho_{K-x}(u)^{\alpha}dx\Big)^{1/\alpha},\quad\quad\alpha\neq 0,$
and
$\log\rho_{{\rm R}_{0}K}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx,\quad\quad\alpha=0.$
For $\alpha\in(-1,0)$, Haddad and Ludwig [25] showed that the radial mean body ${\rm R}_{\alpha}K$ is proportional to $\Pi^{*,-\alpha}\operatorname{1}_{K}$, where $\Pi^{*,-\alpha}\operatorname{1}_{K}$ denotes the fractional polar projection body arsing from the affine fractional Sobolev inequality (1.7) (see Section 2.6 for details). Of particular interest, for $f$ with bounded variation, Haddad and Ludwig [25] proved that
$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|u\|_{\Pi^{*,-\alpha}f}^{-\alpha}=\|u\|_{\Pi^{*}f}.$
In Section 7, we show that ${\rm R}_{\alpha}K$ is proportional to ${\rm L}_{\alpha}\operatorname{1}_{K}$ when $\alpha>n$. Motivated by this relation, we introduce an explicit definition of the functional counterpart ${\rm R}_{\alpha}f$, based on the fractional polar projection body $\Pi^{*,-\alpha}f$ and the star-shaped set ${\rm L}_{\alpha}f$. We note that Haddad and Ludwig [27] also used this notation for radial mean bodies, which in their setting are defined via the fractional $L^{2}$ polar projection body and the star-shaped sets appearing in affine HLS inequalities. Here, we use the same notation ${\rm R}_{\alpha}f$ for an $L^{1}$- extension to keep the notation simple, even though the definition in [27] is slightly different.
In our setting, in particular, for a log-concave function $f\in L^{1}(\mathbb{R}^{n})$, we relate ${\rm R}_{\alpha}f$ to the radial mean bodies of its superlevel sets $\{f\geq r\}$. More precisely, for $\alpha>-1$,
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r),\quad\quad\quad\alpha\neq 0,$
and
$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r),\quad\quad\alpha=0,$
where $d\mu_{f}(r)=\frac{|\{f\geq r\}|}{\|f\|_{1}}dr$.
This unexpected relation naturally leads to the definition of ${\rm R}_{\infty}f$,
${\rm R}_{\infty}f={\rm R}_{\infty}~{\rm supp}(f)={\rm D}~{\rm supp}(f),$
where $f\in L^{1}(\mathbb{R}^{n})$ is a log-concave function with compact support, denoted by ${\rm supp}(f)$. The notation ${\rm D}K$ is the difference body of $K$, defined by ${\rm D}K=\{x-y:x,y\in K\}$. This observation yields the following monotonicity property:
${\rm R}_{\alpha}f\subset{\rm R}_{\beta}f\subset{\rm R}_{\infty}f$
and
$\|f\|_{\infty}\Pi^{*}f\subset\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$
for $1<\alpha<\beta<\infty$. When $f=\operatorname{1}_{K}$, these relations reduce to the geometric case of Gardner and Zhang [18].
Gardner and Zhang also established the reversed inclusion in [18, Theorem 5.5], which requires a more involved argument. For a convex body $K\subset\mathbb{R}^{n}$, they proved
${\rm D}K\subset c_{n,\beta}{\rm R}_{\beta}K\subset c_{n,\alpha}{\rm R}_{\alpha}K\subset n|K|\Pi^{*}K,$ (1.10)
for $1<\alpha<\beta<\infty$. In each inclusion, equality holds if and only if $K$ is a simplex, denoted by $\triangle$. Here $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$, where $B(\alpha+1,n)$ denotes the Beta function.
In (1.10), when $\beta=n$, the left side inclusion implies the Rogers-Shephard inequality [44], which is
$|{\rm D}K|\leq\binom{2n}{n}|K|.$
When $\alpha=n$, the right side inclusion yields Zhang’s inequality [50] (see also Ball [5]):
$n^{-n}\binom{2n}{n}=|\triangle|^{n-1}|\Pi^{*}\triangle|\leq|K|^{n-1}|\Pi^{*}K|,$
which is a reversed form of Petty projection inequality.
To extend these geometric inequalities to their functional analogues, there are two natural ways to associate a convex body with a log-concave function:
$K\mapsto\operatorname{1}_{K}\quad\quad\text{and}\quad\quad K\mapsto e^{-\|\cdot\|_{K}}.$
Using the second correspondence, Alonso-Gutiérrez, Bernués and González Merino [2] extended Zhang’s inequality to log-concave functions, showing that for a log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy\leq 2^{n}n!\|f\|_{1}^{n+1}|\Pi^{*}f|.$ (1.11)
If $f(0)=\|f\|_{\infty}$, equality holds if and only if $f(x)=\|f\|_{\infty}e^{-\|x\|_{\triangle}}$. For further results on Zhang’s inequality and on radial mean bodies, we refer the reader to [3, 4, 20, 21, 24, 28].
In Section 9, we extend (1.10) to both log-concave and $s$-concave functions, which are stated below as the following two theorems.
Theorem 4.
Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f\subset 2\|f\|_{1}\Pi^{*}f$
for $1<\alpha<\beta<\infty$. In each inclusion, there is equality if and only if $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$.
When $\alpha=n$, the inclusion
$\frac{1}{\Gamma(n+1)^{1/n}}{\rm R}_{n}f\subset 2\|f\|_{1}|\Pi^{*}f|,$
recovers (1.11). The inclusion relations in Theorem 4 can be deduced from the results of Alonso-Gutiérrez, Bernués and González Merino [2, Lemma 3.2, Remark 5], but the equality cases were not addressed there. In Section 9, we employ a different approach that allows us to establish Theorem 4 together with a complete characterization of equality.
In proving Theorem 4, we first establish an inclusion relation between ${\rm R}_{\alpha}f$ and ${\rm R}_{\beta}f$, and then consider the limits $\alpha\to-1$ and $\beta\to\infty$ to obtain inclusions between ${\rm R}_{\infty}f$ and $\Pi^{*}f$. However, in this limiting process $\beta\to\infty$ degenerates to the singular set $\{o\}$, so that ${\rm R}_{\infty}f$ does not appear in the general case. This motivates us to consider the embedding $K\mapsto\operatorname{1}_{K}$, leading to the following natural question:
Open problem. For a log-concave function $f\in L^{1}(\mathbb{R}^{n})$ with compact support, does
${\rm R}_{\infty}f\subset c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f\subset n\|f\|_{1}\Pi^{*}f$
hold? Here $c_{n,\alpha}$ is the constant in (1.10).
Unfortunately, this inclusion fails for general log-concave integrable functions with compact support, as we show by a counterexample in Section 9. This failure leads us naturally to consider $s$-concave functions.
Our main result concerns $s$-concave functions $f\in L^{1}(\mathbb{R}^{n})$, where we denote the positive part of $f$ by $f_{+}(x):=\max\{f(x),0\}$. The equality cases in this setting are fully characterized as well.
Theorem 5.
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. For $1<\alpha<\beta<\infty$,
${\rm R}_{\infty}f\subset c_{n,\beta}(s){\rm R}_{\beta}f\subset c_{n,\alpha}(s){\rm R}_{\alpha}f\subset(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$ (1.12)
where $c_{n,\alpha}(s)=(nB(\alpha+1,n+\frac{1}{s}))^{-1/\alpha}$. In each inclusion, there is equality if and only if $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$, where $c>0$, $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.
Theorems 4 and 5 share the same general structure and rely on similar methods of proof. As noted above, in Theorem 4, $\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f$ degenerates to the singular set $\{o\}$ as $\beta\to\infty$, so that the set ${\rm R}_{\infty}f$ does not appear. In contrast, Theorem 5 avoids this degeneration, and in the limit $s\to\infty$, the inclusion (1.12) recovers (1.10).
2. Preliminary
We collect some basic notation and results on $L^{p}$ spaces, Schwarz symmetrals, star-shaped sets, log-concave functions, radial mean bodies and fractional polar projection bodies. The books of Gardner [14], Schneider [45], and Schneider and Weil [46] are good general references.
2.1. $L^{p}$ spaces
Let p>1p>-1 and $(X,\mathcal{B},\mu)$ be a measure space. In this paper, we mainly consider the case $X=\mathbb{R}^{n}$ or $X\subset\mathbb{R}^{n}$ with $\mu$ being the $n$-dimensional Lebesgue measure.
For $p\neq 0$ and a measurable function $f:X\rightarrow\mathbb{R}$, let
$\|f\|_{p}=\Big(\int_{X}|f(x)|^{p}d\mu(x)\Big)^{1/p}.$
In particular, if $p\in(-1,0)$ and $\mu(\{x\in X:f(x)=0\})>0$, we set $\|f\|_{p}=+\infty$. The space $L^{p}(X)$ is then defined as
$L^{p}(X)=\{f:X\rightarrow\mathbb{R}:f~\text{is measurable},~\|f\|_{p}<\infty\},$
where the $f=g$ in $L^{p}(X)$ means that they are equal almost everywhere.
When $\mu(x)=1$, the $L^{p}$-norm is monotone in $p$, that is,
$\|f\|_{p}\leq\|f\|_{q},$
for 1<p<q<0-1<p<q<0 or $0<p<q<\infty$. This follows from Jensen’s inequality. In fact, the monotonicity also holds for $p<0<q$. To see this, under the assumption $\mu(X)=1$, we first note the limit
$\lim_{p\rightarrow 0}\log\|f\|_{p}=\int_{X}\log|f(x)|d\mu(x),$
which follows from the continuity of $\|f\|_{p}$ with respect to $p$. Applying Jensen’s inequality then gives
$\|f\|_{p}\leq\int_{X}\log|f(x)|d\mu(x)\leq\|f\|_{q},$
for $p<0<q$.
2.2. Symmetrization
For $E\subset\mathbb{R}^{n}$, the characteristic function $\operatorname{1}_{E}$ is defined by $\operatorname{1}_{E}(x)=1$ for $x\in E$, and $\operatorname{1}_{E}(x)=0$ for $x\notin E$. If $E$ is a Borel set of finite measure, the Schwarz symmetral of $E$, denoted by $E^{\star}$, is the closed, centered ball with the same volume as $E$.
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a non-negative measurable function. The superlevel set of $f$ is defined by $\{f\geq t\}=\{x\in X:f(x)\geq t\}$ for $t>0$. We say $f$ is non-zero if $\{f\neq 0\}$ has positive measure. For $f$ with superlevel sets of finite measure, by the layer cake formula,
$f(x)=\int_{0}^{\infty}\operatorname{1}_{\{f\geq t\}}(x)dt$
for almost every $x\in\mathbb{R}^{n}$.
The Schwarz symmetral of $f$, denoted by $f^{\star}$, is defined as
$f^{\star}(x)=\int_{0}^{\infty}\operatorname{1}_{\{f\geq t\}^{\star}}(x)dt$
for $x\in\mathbb{R}^{n}$. Hence $f^{\star}$ is radially symmetric and the superlevel set $\{f^{\star}\geq t\}$ has the same volume as $\{f\geq t\}$ for every $t>0$.
One main technique in our proofs is the Riesz rearrangement inequality (see, for example [31, Theorem 3.7]) and the characterization of equality case due to Burchard [8].
Theorem 2.1 (Riesz’s rearrangement inequality.).
For measurable $f,g,k:\mathbb{R}^{n}\rightarrow[0,\infty)$ with superlevel sets of finite measure.
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}f(x)k(x-y)g(y)dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}f^{\star}(x)k^{\star}(x-y)g^{\star}(y)dxdy.$
Theorem 2.2 (Burchard).
Let $A,B$ and $C$ be sets of finite positive measure in $\mathbb{R}^{n}$ and denote by $a,b$ and $c$ the radii of their Schwarz symmetrals $A^{\star},B^{\star}$ and $C^{\star}$. For $|\alpha-\beta|<\gamma<\alpha+\beta$, there is equality in
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{A}(x)\operatorname{1}_{B}(x-y)\operatorname{1}_{C}(y)dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{A^{\star}}(x)\operatorname{1}_{B^{\star}}(x-y)\operatorname{1}_{C^{\star}}(y)dxdy$
if and only if, up to sets of measure zero,
$A=x_{0}+aD,~B=x_{1}+bD,~C=x_{2}+cD,$
where $D$ is a centered ellipsoid, and $x_{0},x_{1}$ and $x_{2}=x_{0}+x_{1}$ are vectors in $\mathbb{R}^{n}$.
2.3. Star-shaped sets and dual mixed volumes
We say $K\subset\mathbb{R}^{n}$ is a star-shaped set (with respect to the origin) if for any $x\in K$ the line segment $[0,x]$ is contained in $K$. The gauge function $\|\cdot\|_{K}:\mathbb{R}^{n}\rightarrow[0,+\infty]$ of a star-shaped set $K$ is defined by
$\|x\|_{K}=\inf\{\lambda>0:x\in\lambda K\},$
[…]
$\rho_{K}(x)=\|x\|_{K}^{-1}=\sup\{\lambda\geq 0:\lambda x\in K\}.$
In particular, when $K$ is the unit ball, $\|\cdot\|_{K}$ is the Euclidean norm $|\cdot|$.
For a star-shaped set $K$ with measurable radial function, the $n$-dimensional volume, or the $n$-dimensional Lebesgue measure of $K$ is given by
$|K|=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n}du,$
where $du$ is the spherical Lebesgue measure. The dual mixed volume for star-shaped sets $K$ and $L$ by Lutwak [34] is given by
$\tilde{V}_{\alpha}(K,L)=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\rho_{L}(u)^{\alpha}du,$
[…]
For $\alpha\in(0,n)$, and star-shaped sets $K$ and $L$ of finite volume, the dual mixed volume inequality states that
$\tilde{V}_{\alpha}(K,L)\leq|K|^{\frac{n-\alpha}{n}}|L|^{\frac{\alpha}{n}},$ (2.1)
where equality holds if and only if $K$ and $L$ are dilates, that is, $\rho_{K}(u)=c\rho_{L}(u)$ for some $c\geq 0$ and almost all $u\in\mathbb{S}^{n-1}$. For $\alpha<0$ and $\alpha>n$, the dual mixed volume inequality states that
$\tilde{V}_{\alpha}(K,L)\geq|K|^{\frac{n-\alpha}{n}}|L|^{\frac{\alpha}{n}},$ (2.2)
where equality holds if and only if $K$ and $L$ are dilates.
2.4. Log-concave functions
A function $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is log-concave if $\log f$ is a concave function on $\mathbb{R}^{n}$, with values in $[-\infty,\infty)$. For $s>0$, a function $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is $s$-concave if $f^{s}$ is concave on its support.
Lemma 2.3.
[2, Lemma 2.1] Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a log-concave and integrable function. Then the function $g:\mathbb{R}^{n}\rightarrow[0,\infty)$ defined by
$g(x)=\int_{\mathbb{R}^{n}}\min\{f(y),f(y+x)\}dy$
is even, log-concave, with $\|g\|_{\infty}=g(o)=\|f\|_{1}$ and ${\rm supp}(g)$ containing the origin point in its interior.
For $\lambda\in[0,1]$, $a,b>0$ and $s>0$, we denote by
[…]
For an $s$-concave function $f$, we have $f((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(f(x),f(y))$ for $\lambda\in[0,1]$. We say $f$ is $(\lambda_{0},s)$-concave, if $f((1-\lambda_{0})x+\lambda_{0}y)\geq\mathcal{M}_{\lambda_{0}}^{s}(f(x),f(y))$ for fixed $\lambda_{0}\in(0,1)$.
The following theorem from Balogh and Kristály [6] gives the equality characterization for Borell-Brascamp-Lieb inequality.
Theorem 2.4.
[6, Theorem 3.1] Let $\lambda\in(0,1)$ and $s>-\frac{1}{n}$. Let $f,g,h:\mathbb{R}^{n}\rightarrow[0,\infty)$ be three non-zero, compactly supported integrable functions satisfying
$h((1-\lambda)x+\lambda y)\geq{\mathcal{M}}_{\lambda}^{s}(f(x),g(y)).$
Then there is equality in
$\int_{\mathbb{R}^{n}}h\geq{\mathcal{M}}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}f,\int_{\mathbb{R}^{n}}g\Big),$
if and only if there is $x_{0}\in\mathbb{R}^{n}$, an $(\lambda_{0},s)$-concave function $\Phi:{\rm supp}(f)\rightarrow[0,\infty)$ with $\lambda_{0}=\frac{\lambda c_{0}}{1-\lambda+\lambda c_{0}}$ and $c_{0}=(\frac{|{\rm supp}(g)|}{|{\rm supp}(f)|})^{1/n}$ such that up to null measure sets,
${\rm supp}(g)=c_{0}\,{\rm supp}(f)+x_{0},~~\text{and}~~{\rm supp}(h)=(1-\lambda+\lambda c_{0})\,{\rm supp}(f)+sx_{0},$
and for almost all $x\in{\rm supp}(f)$,
$\left\{\begin{aligned} &f(x)=\Phi(x)&&,\\ &g(c_{0}x+x_{0})=c_{0}^{\frac{1}{s}}\Phi(x)&&,\\ &h((1-\lambda+\lambda c_{0})x+\lambda x_{0})=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{0}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&.\end{aligned}\right$. (2.3)
2.5. Convex bodies and radial mean bodies
A convex body in $\mathbb{R}^{n}$ is a convex, compact set with non-empty interior. Let $K\subset\mathbb{R}^{n}$ be a convex body. For $\alpha>-1$ and $\alpha\neq 0$, Gardner and Zhang [18] defined the radial $\alpha$-th mean body of $K$ via its radial function by
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{1}{|K|}\int_{K}\rho_{K-x}(u)^{\alpha}dx.$
For $\alpha=0$,
$\log\rho_{{\rm R}_{0}f}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx.$
For $x=x^{\prime}+tu$ with $x^{\prime}\in K|u^{\perp}$, we have
$\int_{K}\rho_{K-x}(u)^{\alpha}dx=\int_{K|u^{\perp}}\int_{t_{1}}^{t_{2}}(t_{2}-t)^{\alpha}dtdx^{\prime},$
where $t_{1}=\min\{t:x^{\prime}+tu\in K\}$ and $t_{2}=\max\{t:x^{\prime}+tu\in K\}$. Since $t_{2}-t_{1}=|K\cap(x^{\prime}+\mathbb{R}u)|$, we then obtain
$\int_{K}\rho_{K-x}(u)^{\alpha}du=\frac{1}{\alpha+1}\int_{K|u^{\perp}}|K\cap(x^{\prime}+\mathbb{R}u)|^{\alpha+1}dx^{\prime}.$
2.6. Fractional polar projection bodies
For $\alpha\in(-1,0)$ and measurable $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$,
$\rho_{\Pi^{*,-\alpha}f}(u)^{-\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x+tu)-f(x)|dxdt.$ (2.4)
Then the affine fractional Sobolev inequality (1.7) can be written as
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|\Pi^{*,-\alpha}f|^{\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy.$
Following the calculation in [27], we obtain that
$\rho_{\Pi^{*,-\alpha}\operatorname{1}_{K}}(u)^{\alpha}$ $=\int_{0}^{\infty}t^{\alpha-1}|K\triangle(K+tu)|dt$ (2.5)
$=\int_{0}^{\infty}t^{\alpha-1}\int_{K|u^{\perp}}2\min\{|K\cap(y+\mathbb{R}u),t|\}dydt$
$=2\int_{K|u^{\perp}}\int_{0}^{\infty}t^{\alpha-1}\min\{|K\cap(y+\mathbb{R}u)|,t\}dtdy$
$=2\int_{K|u^{\perp}}\Big(\int_{0}^{|K\cap(y+\mathbb{R}u)|}t^{\alpha}dt+\int_{|K\cap(y+\mathbb{R}u)|}^{\infty}|K\cap(y+\mathbb{R}u)|t^{\alpha-1}dt\Big)dy$
$=-\frac{2}{\alpha(\alpha+1)}\int_{K|u^{\perp}}|K\cap(y+\mathbb{R}u)|^{\alpha+1}dy,$
where $K\triangle L$ is the symmetric difference of $K,L\subset\mathbb{R}^{n}$. Hence,
$\Pi^{*,-\alpha}\operatorname{1}_{K}=\Big(\frac{2|K|}{-\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K,$ (2.6)
for $\alpha\in(-1,0)$.
3. Proof of Theorem 1
In this section, we focus on the proof of Theorem 1. Before presenting the proof, we first establish the following lemma, which is based on the Minkowski inequality (see, for example, [31, Theorem 2.4]). Let $F$ be a non-negative, measurable function on $\mathbb{R}^{n}\times\mathbb{R}^{n}$ and $p\geq 1$. The Minkowski inequality states that
$\int_{\mathbb{R}^{n}}\Big(\int_{\mathbb{R}^{n}}F(x,y)^{p}dx\Big)^{1/p}dy\geq\Big(\int_{\mathbb{R}^{n}}\Big(\int_{\mathbb{R}^{n}}F(x,y)dy\Big)^{p}dx\Big)^{1/p}.$ (3.1)
Equality and finiteness in (3.1) for $p\in(1,\infty)$ imply the existence of measurable functions $\varphi,\psi:\mathbb{R}^{n}\rightarrow[0,\infty)$ such that $F(x,y)=\varphi(x)\psi(y)$ almost everywhere.
Lemma 3.1.
For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
$\|f\|_{\frac{n}{n+\alpha}}\geq\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
Equality holds if and only if $f=c\operatorname{1}_{E}$ for some measurable $E$ of finite measure and some constant $c>0$.
Proof.
By the Minkowski inequality (3.1),
$\int_{\mathbb{R}^{n}}f(x)^{\frac{n}{n+\alpha}}dx$ $=\int_{\mathbb{R}^{n}}\Big(\int_{0}^{\infty}\operatorname{1}_{\{f\geq r\}}(x)^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}}dx$ (3.2)
$\geq\Big(\int_{0}^{\infty}\Big(\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)dx\Big)^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}}$
$=\Big(\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}},$
that is
$\|f\|_{\frac{n}{n+\alpha}}\geq\int_{0}^{\infty}|\{f>t\}|^{\frac{n+\alpha}{n}}dt.$
It is clear that equality is attained when $f=c\operatorname{1}_{E}$. Conversely, if equality holds in (3.2), there are non-negative measurable functions $\varphi$ and $\psi$ such that
$\operatorname{1}_{\{f\geq r\}}(x)=\varphi(x)\psi(t).$
This implies that
$\{(x,r)\in\mathbb{R}^{n}\times\mathbb{R}:f(x)\geq r\}=\{x\in\mathbb{R}^{n}:\varphi(x)>0\}\times\{r\in\mathbb{R}:\psi(r)>0\}.$
Therefore, for every fixed $r\in\mathbb{R}$ such that $r>0$ and $r\leq f(x)$ for some $x$, the superlevel set $\{f\geq r\}$ is a set independent of $r$, which implies that $f=c\operatorname{1}_{E}$ for some measurable set $E$ of finite measure and some constant $c>0$. ∎
We are now ready to prove Theorem 1.
Proof of Theorem 1.
Note that
$\min\{f(x),f(y)\}=\int_{0}^{\infty}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)dr.$
By Fubini’s theorem and Riesz rearrangement inequality, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy$ $=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)}{|x-y|^{n-\alpha}}dxdydr$
$\leq\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{\{f\geq r\}^{\star}}(y)}{|x-y|^{n-\alpha}}dxdydr$
$=\sigma_{n,\alpha}\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr,$
where the last equation is from the definition of chord power integrals and $|E^{\star}|=|E|$ for any $E\subset\mathbb{R}^{n}$ with finite measure. The best constant $\sigma_{n,\alpha}$ is given by
$\sigma_{n,\alpha}=\frac{nI_{\alpha+1}(B^{n})}{\alpha(\alpha+1)\omega_{n}^{\frac{\alpha}{n}}}.$
Then Lemma 3.1 implies that
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
When the equality holds, it follows from Lemma 3.1 that $f$ is a characteristic function. Then by the isoperimetric inequality of chord integrals, we have $f=c\operatorname{1}_{B}$, up to a set of measure zero. ∎
4. The star-shaped set ${\rm L}_{\alpha}f$
Building on the approach of Haddad and Ludwig [25, 26, 27], we introduce the following anisotropic version of $I_{\alpha+1}(f)$. Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function. We consider
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy,$
where $K\subset\mathbb{R}^{n}$ is a star-shaped set with measurable radial function and $\alpha>0$.
Here, we first recall the definition of ${\rm L}_{\alpha}f$:
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt.$
Using Fubini’s theorem and polar coordinate, we obtain that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy$ $=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(y+z),f(y)\}}{\|z\|_{K}^{n-\alpha}}dzdy$
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\rho_{K}(tu)^{n-\alpha}t^{n-1}\min\{f(y),f(y+tu)\}dtdudy$
$=\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(y),f(y+tu)\}dydtdu$
$=\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}du.$
That is,
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f),$ (4.1)
for measurable $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ and star-shaped set $K\subset\mathbb{R}^{n}$ with measurable radial function. In particular,
$|{\rm L}_{n}f|$ $=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{{\rm L}_{n}f}(u)^{n}du$
$=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}t^{n-1}\min\{f(x),f(x+tu)\}dxdtdu$
$=\frac{1}{n}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy.$
Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function. Then the following hold:
For $\alpha\in(0,n)$, the dual mixed volume inequality (2.1) together with (4.1) implies
$\sup\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
For $\alpha>n$, the dual mixed volume inequality (2.2) together with (4.1) yields
$\inf\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
Moreover, in either regime a suitable dilate of ${\rm L}_{\alpha}f$ attains the supremum (respectively, infimum), provided $|{\rm L}_{\alpha}f|$ is finite.
We conclude this section with a result on the convexity of ${\rm L}_{\alpha}f$ when $f$ is log-concave.
Proposition 4.1.
If $f:\mathbb{R}^{n}\to[0,\infty)$ is log-concave and in $L^{1}(\mathbb{R}^{n})$, then ${\rm L}_{\alpha}f$ is a convex set for every $\alpha>0$.
Proof.
For $u\in\mathbb{S}^{n-1}$, we define
$g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx.$
Since $f$ is log-concave, Lemma 2.3 implies that $g$ is also log-concave.
Note that
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}g(tu)dt.$
Applying a result of Ball [5, Theorem 5], we obtain that
$u\mapsto\left(\int_{0}^{\infty}t^{\alpha-1}g(tu)dt\right)^{-\frac{1}{\alpha}}$
defines a norm on $\mathbb{R}^{n}$, which concludes the proof. ∎
5. Affine isoperimetric inequalities for $\alpha\in(0,n)$
In this section, we aim to establish an affine version of Theorem 1. To this end, we first prove the following crucial lemma, which uses the Riesz rearrangement inequality. We say that $A\subset\mathbb{R}^{n}$ is homothetic to $B\subset\mathbb{R}^{n}$ if there exist $x\in\mathbb{R}^{n}$ and $a>0$ such that $A=x+aB$.
Lemma 5.1.
Let $q>0$ and $K\subset\mathbb{R}^{n}$ be a star-shaped set with measurable radial function and $|K|>0$. For a measurable function $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ such that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy<\infty,$
there is
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{\|x-y\|_{K^{\star}}^{q}}dxdy.$
The equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost all $r>0$, the level set $\{f\geq r\}$ has measure zero or is homothetic to $D$ up to a set of measure zero.
Proof.
[…]
$\|z\|_{K}^{-q}=\int_{0}^{\infty}\operatorname{1}_{t^{-1/q}K}(z)dt.$
Using the Fubini’s theorem, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy$ $=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)}{\|x-y\|_{K}^{q}}dxdydr$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}(y)dxdydtdr.$
By the Riesz rearrangement inequality, Theorem 2.2, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}$ $(y)dxdy$
$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy$
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f^{\star}\geq r\}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f^{\star}\geq r\}}(y)dxdy,$
which implies that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{\|x-y\|_{K^{\star}}^{q}}dxdy.$
If the equality holds, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}$ $(y)dxdy$
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy$
for almost all $(r,t)\in(0,\infty)^{2}$. For sufficiently large $t>0$, the assumptions of Theorem 2.2 are fulfilled. Therefore, up to sets of measure zero, we have
$\{f\geq r\}=x+aD,~~t^{-1/q}K=bD,$
where $D$ is an ellipsoid and $x\in\mathbb{R}^{n}$. Since $K=(\frac{|K|}{|D|})^{1/n}D$, the ellipsoid $D$ is not dependent on $t$ and $r$, which concludes the proof. ∎
Lemma 5.2.
Let $\alpha\in(0,n)$. For non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
$|{\rm L}_{\alpha}f|\leq|{\rm L}_{\alpha}f^{\star}|.$
If $|{\rm L}_{\alpha}f^{\star}|<\infty$, there is equality if and only if $f$ is a translate of $f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi$.
Proof.
We first assume that $|{\rm L}_{\alpha}f|<\infty$. By letting $q=n-\alpha$ in Lemma 5.1 and combining it with (4.1), we obtain
$\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f)\leq\tilde{V}_{\alpha}(K^{\star},{\rm L}_{\alpha}f^{\star}).$
With $K={\rm L}_{\alpha}f$ and the dual mixed volume inequality (2.1), we obtain
$|{\rm L}_{\alpha}f|$ $=\tilde{V}_{\alpha}({\rm L}_{\alpha}f,{\rm L}_{\alpha}f)$
$\leq\tilde{V}_{\alpha}(({\rm L}_{\alpha}f)^{\star},{\rm L}_{\alpha}f^{\star})$
$\leq|({\rm L}_{\alpha}f)^{\star}|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}$
$=|{\rm L}_{\alpha}f|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}},$
which implies that
$|{\rm L}_{\alpha}f|\leq|{\rm L}_{\alpha}f^{\star}|.$
The equality case follows from Lemma 5.1.
If $|{\rm L}_{\alpha}f|=\infty$, define
$f_{k}(x)=f(x)\operatorname{1}_{kB^{n}}(x),~~k\in\mathbb{N}.$
By the monotone convergence theorem,
$\lim_{k\rightarrow\infty}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f_{k}(x),f_{k}(x+tu)\}dxdt=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,$
with the convergence being monotone increasing. Applying the monotone convergence theorem once more then gives
$\lim_{k\rightarrow\infty}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{$ $f_{k}(x),f_{k}(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}du$
$=\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}du.$
That is,
$\lim_{k\rightarrow\infty}|{\rm L}_{\alpha}f_{k}|=|{\rm L}_{\alpha}f|=\infty.$ (5.1)
Since $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$ and $f_{k}$ has compact support, $|{\rm L}_{\alpha}f_{k}|<\infty$. By the definition of the Schwarz symmetral, we have $(f_{k})^{\star}\leq f^{\star}$. Then the first part of the proof implies that
$|{\rm L}_{\alpha}f_{k}|\leq|{\rm L}_{\alpha}(f_{k})^{\star}|\leq|{\rm L}_{\alpha}f^{\star}|.$
Taking the limit as $k\rightarrow\infty$ in the previous inequality, we obtain from (5.1) that
$|{\rm L}_{\alpha}f^{\star}|=\infty,$
which gives the desired result. ∎
Proof of Theorem 2.
By (4.1) and the dual mixed volume inequality (2.1), we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f)\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
There is equality precisely when ${\rm L}_{\alpha}f$ is a ball, which is the case for radially symmetric functions.
For the second inequality, noting that $\|f\|_{\frac{n}{n+\alpha}}=\|f^{\star}\|_{\frac{n}{n+\alpha}}$ and ${\rm L}_{\alpha}f^{\star}$ is a ball, we obtain from Theorem Theorem 1, (4.1), and Lemma 5.2 that
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}$ $\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy$
$=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f^{\star})$
$=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}$
$\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
If equality holds throughout, it follows from Theorem Theorem 1 that $f=c\operatorname{1}_{B}$ with $B$ a ball. By Lemma 5.2, $f=f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. This characterizes the equality cases. ∎
6. Affine isoperimetric inequalities for $\alpha>n$
As mentioned in the introduction, Theorem 1 extends the classical isoperimetric inequality for chord power integrals to function. It also provides an $L^{1}$ analogue of the HLS inequalities, for which both inequalities reverse when $\alpha>n$. In our setting, however, this reversal does not appear in such a simple form here.
Before addressing this nontrivial case, we first present the following lemma, which serves as a counterpart to Lemma 5.1
Lemma 6.1.
Let $q>0$ and $K\subset\mathbb{R}^{n}$ be a star-shaped set with measurable radial function and $|K|>0$. If $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is an integrable function such that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}\|x-y\|_{K}^{q}dxdy<\infty,$
then
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}\|x-y\|_{K}^{q}dxdy\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f^{\star}(x),f^{\star}(y)\}\|x-y\|_{K^{\star}}^{q}dxdy.$
Equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost $r>0$, the level set $\{f\geq r\}$ has measure zero or is homothetic to $D$ up to sets of measure zero.
Proof.
[…]
$\|z\|_{K}^{q}=\int_{0}^{\infty}\operatorname{1}_{\mathbb{R}^{n}\setminus t^{1/q}K}(z)dt.$
Using Fubini’s theorem, we obtain
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),$ $f(y)\}\|x-y\|_{K}^{q}dxdy$ (6.1)
$=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\|x-y\|_{K}^{q}\operatorname{1}_{\{f\geq r\}}(y)dxdydr$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\mathbb{R}^{n}\setminus t^{1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}(y)dxdydrdt$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)(1-\operatorname{1}_{t^{1/q}K}(x-y))\operatorname{1}_{\{f\geq r\}}(y)dxdydrdt.$
The Riesz rearrangement inequality, Theorem 2.2, implies that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{1/q}K}(x-y)\operatorname{1}$ ${}_{\{f\geq r\}}(y)dxdy$ (6.2)
$\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy.$
Note that $\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)dx<\infty$ for almost all $r>0$, as $f\in L^{1}(\mathbb{R}^{n})$. Combing (6.2) with the calculation in (6.1), we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}\|x-y\|_{K}^{q}dxdy\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f^{\star}(x),f^{\star}(y)\}\|x-y\|_{K^{\star}}^{q}dxdy.$ (6.3)
Moreover, if equality holds in (6.3), then there exists a null set $N\subset(0,\infty)^{2}$ such that equality holds in (6.2) for all $(r,t)\in(0,\infty)^{2}\setminus N$.
For almost every $t\in(0,\infty)$, we have $(r,t)\in(0,\infty)^{2}\setminus N$ for almost every $r>0$. For $t$ sufficiently small, Theorem 2.2 guarantees the existence of a centered ellipsoid $D$ and $x\in\mathbb{R}^{n}$ such that
$\{f\geq r\}=x+aD\quad\text{and }\quad t^{1/q}K=bD.$
Since $K=(\frac{|K|}{|D|})^{1/n}D$, the ellipsoid $D$ is independent of $r$ and $t$, which concludes the proof. ∎
With the previous lemma established, for $\alpha>n$, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy$ $\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy$ (6.4)
$=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f^{\star}\geq r\}}(x)\operatorname{1}_{\{f^{\star}\geq r\}}(y)}{|x-y|^{n-\alpha}}dxdy$
$=\sigma_{n,\alpha}\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
However, Lemma 3.2 is not applicable in this case because the inequality holds in the opposite direction. We therefore apply a different approach to obtain the result as follows.
Theorem 6.2.
Let $\alpha>n$ and $f$ be a non-zero, non-negative function in $L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$. Then
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
Equality holds if and only if $f=c\operatorname{1}_{B}$ with $B$ a ball and $c>0$.
Proof.
As we calculated in (6.4),
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy\geq\sigma_{n,\alpha}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
By Jensen’s inequality, we have
$\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr\geq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|dr\Big)^{\frac{n+\alpha}{n}}=\|f\|_{\infty}^{-\frac{n+\alpha}{n}}\|f\|_{1}^{\frac{n+\alpha}{n}},$ (6.5)
and thus
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$ (6.6)
If equality holds in (6.6), then equality must occur in the application of Jensen’s inequality in (6.5). This in turn implies that $|\{f\geq r\}|$ is constant for almost every $r\in(0,\|f\|_{\infty})$. Equivalently, this happens if and only if $f=c\operatorname{1}_{E}$, where $E$ is a measurable set of finite measure and $c>0$.
When $f=c\operatorname{1}_{E}$, the inequality (6.6) reduces to the geometric case:
$I_{\alpha+1}(E)\geq\Big(\frac{I_{\alpha+1}(B^{n})}{\omega_{n}^{(n+\alpha)/n}}\Big)|E|^{\frac{n+\alpha}{n}},$
with equality if and only if $E$ is a ball. This then completes the characterization of the equality cases. ∎
Lemma 6.3.
Let $\alpha>n$ and $f:\mathbb{R}^{n}\to[0,\infty)$ be integrable. If $|{\rm L}_{\alpha}f|<\infty$, then
$|{\rm L}_{\alpha}f|\geq|{\rm L}_{\alpha}f^{\star}|.$
Equality holds if and only if $f$ is a translate of $f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
Proof.
By letting $q=n-\alpha$ in Lemma 6.2 and combining it with (4.1), we obtain
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f)\geq n\tilde{V}_{\alpha}(K^{\star},{\rm L}_{\alpha}f^{\star}).$
With $K={\rm L}_{\alpha}f$ and the dual mixed volume inequality (2.2), we have
$|{\rm L}_{\alpha}f|$ $\geq\tilde{V}_{\alpha}({\rm L}_{\alpha}f,{\rm L}_{\alpha}f)$
$\geq\tilde{V}_{\alpha}(({\rm L}_{\alpha}f)^{\star},{\rm L}_{\alpha}f^{\star})$
$\geq|({\rm L}_{\alpha}f)^{\star}|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}$
$=|{\rm L}_{\alpha}f|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}},$
which implies that
$|{\rm L}_{\alpha}f|\geq|{\rm L}_{\alpha}f^{\star}|.$
The equality cases follows from Lemma 6.1. ∎
Building on the results established above, we proceed to the proof of Theorem 3.
Proof of Theorem 3.
For the first inequality, if the left side is $\infty$, the inequality holds trivially. We then assume that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy<\infty.$
It follows from (4.1) and the dual mixed volume inequality (2.2) that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f)\geq n\omega^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
By the equality case in the dual mixed volume inequality, ${\rm L}_{\alpha}f$ must be a ball, which occurs when $f$ is radially symmetric.
For the second inequality, we may assume that $|{\rm L}_{\alpha}f|$ is finite. Note that ${\rm L}_{\alpha}f^{\star}$ is a ball. Then Lemma 6.3 and (4.1) implies that
$n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f^{\star})=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy.$
Combining Theorem 6.2 for the right-hand side with the facts $\|f\|_{1}=\|f^{\star}\|_{1}$ and $\|f\|_{\infty}=\|f^{\star}\|_{\infty}$, we deduce
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
The equality characterization is given by Theorem 6.2 and Lemma 6.3. ∎
7. Radial mean bodies
In this section, we begin by recalling some basic notation and properties of the classical radial mean bodies. Let $K\subset\mathbb{R}^{n}$ be a convex body. For $\alpha>-1$ and $u\in\mathbb{S}^{n-1}$, Gardner and Zhang [18] defined the radial $\alpha$-th mean body of $K$ by its radial function:
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{1}{|K|}\int_{K}\rho_{K-x}(u)^{\alpha}dx$
for $\alpha\neq 0$ and by
$\log\rho_{{\rm R}_{0}K}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx$
for $\alpha=0$.
They further defined ${\rm R}_{\infty}K$ consistently by
$\rho_{{\rm R}_{\infty}K}(u)=\lim_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}K}(u)=\max_{x\in K}\rho_{K-x}(u),$
so that ${\rm R}_{\infty}K$ coincides with the difference body of $K$, namely
${\rm R}_{\infty}K=DK=K-K.$
As introduced in Section 2, $\rho_{{\rm R}_{\alpha}K}(u)$ can be expressed as the integral of the chord length $|K\cap(x^{\prime}+\mathbb{R}u)|$. It is also closely related to the covariogram of $K$, that is
$g_{K}(x)=|K\cap(K+x)|.$
In fact
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{\alpha}{|K|}\int_{0}^{\rho_{DK}(u)}t^{\alpha-1}g_{K}(tu)dt$
for $\alpha>0$.
Note that, for $\alpha>0$,
$\rho_{{\rm L}_{\alpha}\operatorname{1}_{K}}(u)^{\alpha}$ $=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{\operatorname{1}_{K}(x),\operatorname{1}_{K}(x+tu)\}dxdt$
$=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\operatorname{1}_{K}(x)\operatorname{1}_{K}(x+tu)dxdt$
$=\int_{0}^{\infty}t^{\alpha-1}|K\cap(K+tu)|dt$
$=\int_{0}^{\infty}t^{\alpha-1}g_{K}(tu)dt.$
Hence,
${\rm L}_{\alpha}\operatorname{1}_{K}=\Big(\frac{|K|}{\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K$ (7.1)
for $\alpha>0$. Recall that in Section 2.5,
$\Pi^{*,-\alpha}\operatorname{1}_{K}=\Big(\frac{2|K|}{-\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K$ (7.2)
for $\alpha\in(-1,0)$.
Combining (7.1) and (7.2) with Theorem 2 and Theorem 3, we obtain the following Pólya–Szegö inequality:
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\leq\Big(\frac{|{\rm R}_{\alpha}K|}{|{\rm R}_{\alpha}{B^{n}}|}\Big)^{\alpha/n}\leq\Big(\frac{|K|}{|B^{n}|}\Big)^{(n+\alpha)/n},$
for $\alpha\in(0,n)$ and
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\geq\Big(\frac{|{\rm R}_{\alpha}K|}{|{\rm R}_{\alpha}{B^{n}}|}\Big)^{\alpha/n}\geq\Big(\frac{|K|}{|B^{n}|}\Big)^{(n+\alpha)/n},$
for $\alpha>n$. These two formulas can also be regarded as the affine isoperimetric inequalities for chord power integrals.
8. Radial mean bodies for log-concave functions
By (7.1) and (7.2), we define the radial mean bodies for log-concave functions, by
${\rm R}_{\alpha}f=\Big(\frac{-\alpha}{2\|f\|_{1}}\Big)^{1/\alpha}\Pi^{*,-\alpha}f$ (8.1)
for $\alpha\in(-1,0)$ and
${\rm R}_{\alpha}f=\Big(\frac{\alpha}{\|f\|_{1}}\Big)^{1/\alpha}{\rm L}_{\alpha}f$
for $\alpha>0$.
The following proposition describes the relation between ${\rm R}_{\alpha}f$ and the radial $\alpha$-th mean body of the superlevel set $\{f\geq r\}$, for $\alpha\in(-1,0)\cup(0,\infty)$.
Proposition 8.1.
Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave. For $\alpha\in(-1,0)\cup(0,\infty)$,
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r),$
where $d\mu_{f}(r)=\frac{|\{f\geq r\}|}{\|f\|_{1}}dr$.
Proof.
First we consider $\alpha>0$. Recall that
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{\alpha}{\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,$
where
$\min\{f(x),f(x+tu)\}=\int_{0}^{\|f\|_{\infty}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(x+tu)dr.$
By Fubini’s theorem, we then obtain
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{\alpha}{\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\cap(\{f\geq r\}+tu)|dtdr.$
Note that for a convex body $K$,
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{\alpha}{|K|}\int_{0}^{\infty}t^{\alpha-1}|K\cap(K+tu)|dt.$
We have
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}\cdot\frac{|\{f\geq r\}|}{\|f\|_{1}}dr=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r).$
Second, when $\alpha\in(-1,0)$, the definition (8.1) and (2.4) imply that
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x+tu)-f(x)|dxdt.$
Since
$|f(x+tu)-f(x)|=\int_{0}^{\infty}|\operatorname{1}_{\{f\geq r\}}(x+tu)-\operatorname{1}_{\{f\geq r\}}(x)|dr,$
we obtain
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}$ $=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\int_{0}^{\|f\|_{\infty}}|\operatorname{1}_{\{f\geq r\}}(x+tu)-\operatorname{1}_{\{f\geq r\}}(x)|drdxdt$
$=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\triangle(\{f\geq r\}+tu)|dtdr.$
By the calculation in (2.5) and (2.6), we have
$\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}=\frac{-\alpha}{2\{f\geq r\}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\triangle(\{f\geq r\}+tu)|dt,$
which implies
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\frac{2|\{f\geq r\}|}{-\alpha}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r).$
This completes the proof. ∎
By the continuity of the $L^{p}$-norm in $p$, we define
$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r).$ (8.2)
A more explicit formula for ${\rm R}_{0}f$ will be presented in Section 9.
We recall the following convergence result from [25], which shows that as $\alpha\rightarrow-1$, the radial mean body ${\rm R}_{\alpha}f$ converges to the polar projection body of $f$. The opposite endpoint $\alpha=\infty$ will be considered below.
Lemma 8.2.
[25, Theorem 3] For $\alpha\in(-1,0)$ and log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|f\|_{1}\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\rho_{\Pi^{*}f}(u)^{-1}$
for $u\in\mathbb{S}^{n-1}$.
Lemma 8.3.
Let $f\in L^{1}(\mathbb{R}^{n})$ be log-concave and non-zero. If $f$ has compact support, then
$\lim_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)=\sup_{r\in(0,1)}\rho_{{\rm D}\{f\geq r\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u),$
for $u\in\mathbb{S}^{n-1}$.
Before proving the lemma below, we recall a result from [18, Theorem 5.5], which states that for every convex body $K$
${\rm D}K\subset c_{n,\alpha}{\rm R}_{\alpha}K,$
where $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$. The complete statement of the theorem will be given in Section 9.
Proof.
Since Jensen’s inequality implies that ${\rm R}_{\alpha}K\subset{\rm D}K$ for $\alpha>-1$, we have
$\rho_{{\rm R}_{\alpha}f}(u)\leq\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r)\Big)^{1/\alpha}.$
Letting $\alpha\rightarrow\infty$, we obtain
$\limsup_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\leq\sup_{r\in(0,1)}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u).$ (8.3)
On the other hand, by the reverse inclusion ${\rm D}K\subset c_{n,\alpha}{\rm R}_{\alpha}K$,
$\rho_{{\rm R}_{\alpha}f}(u)\geq c_{n,\alpha}^{-1}\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f>r\}}(u)^{\alpha}d\mu_{f}(r)\Big)^{1/\alpha},$
where $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$. Note that $\lim_{\alpha\rightarrow\infty}c_{n,\alpha}=1$. Therefore we have
$\liminf_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\geq\sup_{r}\rho_{{\rm D}\{f\geq r\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u).$ (8.4)
Combining (8.3) with (8.4), we then conclude the proof. ∎
Definition 8.4.
Let $f\in L^{1}(\mathbb{R}^{n})$ be log-concave with compact support. The body ${\rm R}_{\infty}f$ is given by
${\rm R}_{\infty}f={\rm D}{\rm supp}(f).$
If $f$ is not compactly supported, ${\rm R}_{\infty}f=\mathbb{R}^{n}$.
Recall that in the geometric case, Jensen’s inequality implies that
${\rm R}_{\alpha}K\subset{\rm R}_{\beta}K\subset{\rm D}K,$ (8.5)
where Gardner and Zhang [18] further showed that
$\Pi^{*}K\subset((1+\alpha)|K|)^{1/\alpha}{\rm R}_{\alpha}K\subset((1+\beta)|K|)^{1/\beta}{\rm R}_{\beta}K\subset{\rm D}K$ (8.6)
for $1<\alpha<\beta$.
Applying Jensen’s inequality together with (8.5), we immediately obtain
${\rm R}_{\alpha}f\subset{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$
for $1<\alpha<\beta$. The following theorem provides a functional analogue of (8.6).
Theorem 8.5.
For non-zero, log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\|f\|_{\infty}\Pi^{*}f\subset\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$
for $1<\alpha<\beta$.
Proof.
If $f$ is not compactly supported, the right inclusion is trivial since ${\rm R}_{\infty}f=\mathbb{R}^{n}$. We now assume that $f$ is with compact support.
By Proposition 8.1,
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}(1+\alpha)|\{f\geq r\}|\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha}$
$=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\alpha,r}{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha},$
where $l_{\alpha,r}=((1+\alpha)|\{f\geq r\}|)^{1/\alpha}$ It follows from (8.6) that
$l_{\alpha,r}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)\leq l_{\beta,r}\rho_{{\rm R}_{\beta}\{f\geq r\}}(u),$
for $1<\alpha<\beta$.
Jensen’s inequality then implies that
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,r}{\rm R}_{\beta}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha}$
$\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,r}{\rm R}_{\beta}\{f\geq r\}}(u)^{\beta}dr\Big)^{1/\beta}$
$=\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}\rho_{{\rm R}_{\beta}f}(u),$
which gives
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f.$
Letting $\beta\rightarrow\infty$, together with $\lim_{\beta\rightharpoonup\infty}(1+\beta)^{1/\beta}=1$, we obtain the right inclusion. Letting $\alpha\rightarrow-1$, by Lemma 8.2, we obtain the left inclusion. ∎
9. Reversed inequalities
In the previous section, we recalled the monotonicity relations (8.5) and (8.6) for radial mean bodies and extended them to their functional analogues. In this section, we turn to the reverse inclusion, which is considerably more delicate.
We first recall the following theorem established by Gardner and Zhang [18]. In particular, for $\alpha=n$,
$c_{n,n}{\rm R}_{n}K\subset n|K|\Pi^{*}K$
which yields Zhang’s inequality [50].
Theorem 9.1.
[18, Theorem 5.5] Let $K$ be a convex body in $\mathbb{R}^{n}$. If $1<\alpha<\beta$, then
${\rm D}K\subset c_{n,\beta}{\rm R}_{\beta}K\subset c_{n,\alpha}{\rm R}_{\alpha}K\subset n|K|\Pi^{*}K.$
In each inclusion equality holds if and only if $K$ is a simplex.
To extend Theorem 9.1 to the functional setting, we rely on the following result from [27]. It generalizes the classical result of Marshall, Olkin and Proschan [41], whose simple proof was later given by Milman and Pajor [43].
Lemma 9.2 ([27], Lemma 17).
Let $\omega:[0,\infty)\to[0,\infty)$ be decreasing with
$0<\int_{0}^{\infty}t^{\alpha-1}\omega(t)dt<\infty$
for $\alpha>0$ and
$0<\int_{0}^{\infty}t^{\alpha-1}(\omega(0)-\omega(t))dt<\infty$
for $\alpha\in(-1,0)$. If $\varphi:[0,\infty)\to[0,\infty)$ is non-zero, with $\varphi(0)=0$, and such that $t\mapsto\varphi(t)$ and $t\mapsto{\varphi(t)}/{t}$ are increasing on $(0,\infty)$, then
$\zeta(\alpha)=\left\{\begin{aligned} &\left(\frac{\int_{0}^{\infty}t^{\alpha-1}\omega(\varphi(t))dt}{\int_{0}^{\infty}t^{\alpha-1}\omega(t)dt}\right)^{1/\alpha}&&\text{for}~~\alpha>0,\\ &\exp\left(\int_{0}^{\infty}\frac{\omega(\varphi(t))-\omega(t)}{t\omega(0)}dt\right)&&\text{for}~~\alpha=0,\\ &\left(\frac{\int_{0}^{\infty}t^{\alpha-1}(\omega(\varphi(t))-\omega(0))dt}{\int_{0}^{\alpha-1}(\omega(t)-\omega(0))dt}\right)^{1/\alpha}&&\text{for}~~-1<\alpha<0.\end{aligned}\right$.
is a continuous, decreasing function of $\alpha$ on $(-1,\infty)$. Moreover, $\zeta$ is constant on $(-1,\infty)$ if $\varphi(t)=\lambda t$ on $0,\infty)$ for some $\lambda>0$.
In particular, taking $\omega(t)=e^{-t}$ and $\omega(t)=(1-st)_{+}^{1/s}$ with $s>0$ yields the classical analytic continuation formulas for the Gamma and Beta functions,
[…]
and
[…]
We note here, in the setting of Lemma 9.2, that if $\zeta(\alpha)$ is constant on $(-1,\infty)$, say $\zeta(\alpha)=\lambda$, then $\varphi(t)=\lambda t$ for all $t\geq 0$. When $\omega(t)$ is strictly decreasing, this conclusion follows from the proof of Milman and Pajor, but for completeness, we include a proof below.
Lemma 9.3.
Let $\omega$, $\varphi$, and $\zeta$ be as in Lemma 9.2. Suppose $\omega$ is strictly decreasing and $1<\alpha<\beta<\infty$. Then $\zeta(\alpha)=\zeta(\beta)$ implies $\varphi(t)=\lambda t$ with $\lambda=\zeta(\alpha)^{-1}$.
Proof.
Let $\lambda=\zeta(\alpha)^{-1}$. We first consider $0<\alpha<\beta<\infty$. For $\alpha>0$, we then have
$\int_{0}^{\infty}\omega(\lambda t)t^{\alpha-1}dt=\int_{0}^{\infty}\omega(\varphi(t))t^{\alpha-1}dt.$
We set $\Psi(r)=\int_{r}^{\infty}(\omega(\lambda t)-\omega(\varphi(t)))dt$, where $\Psi(0)=\Psi(\infty)=0$. Since $\varphi(t)/t$ is increasing while $\omega$ is decreasing, we have
$\omega(\varphi(t))>\omega(\lambda t),~~\text{if}~~{\varphi(t)}/{t}<\lambda,$
and
$\omega(\varphi(t))\leq\omega(\lambda t),~~\text{if}~~{\varphi(t)}/{t}\geq\lambda.$
This implies the sign of $\omega(\lambda r)-\omega(\varphi(r))=-\Psi^{\prime}(r)$. Therefore, $\Psi$ is first increasing and then decreasing, which yields that $\Psi(r)\geq 0$ for all $r\geq 0$. That is
$\int_{r}^{\infty}\omega(\lambda t)t^{\alpha-1}dt\geq\int_{r}^{\infty}\omega(\varphi(t))dt.$
For $\beta>\alpha>0$, we have
$\int_{0}^{\infty}\omega(\varphi(t))t^{\beta-1}dt$ $=(\beta-\alpha)\int_{0}^{\infty}r^{\beta-\alpha-1}\Big(\int_{r}^{\infty}\omega(\varphi(t))r^{\alpha-1}dt\Big)dr$
$\leq(\beta-\alpha)\int_{0}^{\infty}r^{\beta-\alpha-1}\Big(\int_{r}^{\infty}\omega(\lambda t)t^{\alpha-1}dt\Big)dr$
$=\int_{0}^{\infty}\omega(\lambda t)t^{\beta-1}dt=\frac{1}{\lambda^{\beta}}\int_{0}^{\infty}\omega(t)t^{\beta-1}dt,$
which implies $\zeta(\alpha)\geq\zeta(\beta)$. As we assumed, $\zeta(\alpha)=\zeta(\beta)=\lambda^{-1}$, equality holds in the inequality above, which implies that $\Psi(r)=0$ for all $r\geq 0$. Then we have
$\Psi^{\prime}(r)=\omega(\varphi(t))-\omega(\lambda t)=0.$
Since $\omega$ is strictly decreasing, we obtain $\varphi(t)=\lambda t$.
For $1<\alpha<\beta<0$, the proof is the same. Then we consider $1<\alpha<0<\beta<\infty$. Since $\zeta$ is decreasing, we have
$\zeta(\beta)\leq\zeta(0)\leq\zeta(\alpha).$
As assumed $\zeta(\alpha)=\zeta(\beta)=\lambda^{-1}$, we obtain that $\zeta$ is a constant on $(0,\beta)$, by the first case, we have $\varphi(t)=\lambda t$. ∎
9.1. Results for log-concave functions
We begin this subsection by recalling a result from [2], which will serve as a main tool in the equality characterization in Theorem 3.
Lemma 9.4.
[2, Lemma 4.2] Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave such that $f(o)=\|f\|_{\infty}$. Let
$g(x)=\int_{\mathbb{R}^{n}}\min\{f(y),f(y+x)\}dy.$
Then for every $x\in\mathbb{R}^{n}$ and $\lambda\in[0,1]$, $g(\lambda x)=g(o)^{1-\lambda}g(x)^{\lambda}$ if and only if $f(x)=\|f\|_{\infty}e^{-\|x\|_{\triangle}}$ with $\triangle$ a simplex containing the origin.
For the reader’s convenience, we restate Theorem 4 below before giving the proof.
Theorem 9.5.
Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave. Then
$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f\subset 2\|f\|_{1}\Pi^{*}f$ (9.2)
for $1<\alpha<\beta<\infty$. In each inclusion, equality holds if and only if $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$, where $\triangle$ a simplex containing the origin.
Proof.
By Lemma 2.3,
$g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
is an even, log-concave function and attains maximum at $t=0$.
For fixed $u\in\mathbb{S}^{n-1}$, we apply Lemma 9.2 to $\omega(t)=g(0)e^{-t}$ and $\varphi(t)=-\log(g(tu)/g(0))$. We observe that
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}g(tu)dt}{\int_{0}^{\infty}t^{\alpha-1}g(0)e^{-t}dt}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
for $\alpha>0$.
For $\alpha\in(-1,0)$, note that
$g(0)-g(tu)$ $=\|f\|_{1}-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ (9.3)
$=\frac{1}{2}\int_{\mathbb{R}^{n}}f(x)dx+\frac{1}{2}\int_{rn}f(x+tu)dx-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
$=\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x)-f(x+tu))_{+}dx+\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x+tu)-f(x))_{+}dx$
$=\frac{1}{2}\int_{\mathbb{R}^{n}}|f(x)-f(x+tu)|dx,$
where $h(x)_{+}$ denotes the positive part of the function $h(x)$. Then we have
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}(g(0)-g(tu))dt}{\int_{0}^{\infty}t^{\alpha-1}g(0)(1-e^{-t})dt}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
for $\alpha\in(-1,0)$.
Therefore Lemma 9.2 implies that
$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f.$ (9.4)
Letting $\alpha\rightarrow-1$, Lemma 8.2 gives the right inclusion. Since $\omega(t)=e^{-t}$ is strictly decreasing, by Lemma 9.3, equality holds in (9.4) if and only if $g(tu)=g(0)e^{-\lambda t}$ for some $\lambda>0$. Then Lemma 9.4 shows that $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$, where $\triangle$ is a simplex containing the origin. ∎
In fact, if we set $f(x)=e^{-\|x\|_{K}}$ for a convex body $K\subset\mathbb{R}^{n}$, then inclusion (9.2) in Theorem 9.5 reduces to the geometric case stated in Theorem 9.1. However, in this functional extension, the endpoint ${\rm R}_{\infty}f$ is $\mathbb{R}^{n}$, so the inclusion no longer extends to that limit case.
Furthermore, combining Lemma 9.2 with (9.4), we obtain an explicit expression for $\zeta(0)$ by the continuity of $\zeta(\alpha)$, where
$\log\zeta(0)=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx-e^{-t}\Big)dt.$ (9.5)
When $f=\operatorname{1}_{K}$, the quantity $\log\zeta(0)=\log\rho_{{\rm R}_{0}K}(u)$. The next result confirms that the formula (9.5) is consistent with the definition of ${\rm R}_{0}f$ (8.2), namely $\rho_{{\rm R}_{0}f}(u)=\zeta(0)$ with $u\in\mathbb{S}^{n-1}$ fixed.
Proposition 9.6.
For fixed $u\in\mathbb{S}^{n-1}$ and non-zero, log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\rho_{{\rm R}_{0}f}(u)=\zeta(0).$
Proof.
By (9.5),
$\log\zeta(0)=$ $-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx-e^{-t}\Big)dt$
$=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\int_{0}^{\|f\|_{\infty}}|\frac{\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}d\mu_{f}(r)-e^{-t}\Big)dt$
$=\int_{0}^{\|f\|_{\infty}}\Big(-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{|\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}-e^{-t}\Big)dt\Big)d\mu_{f}(r).$
Note that
$\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)$ $=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{|\{f\geq r\}|}\int_{\mathbb{R}^{n}}\min\{\operatorname{1}_{\{f\geq r\}}(x),\operatorname{1}_{\{f\geq r\}}(x+tu)\}dx-e^{-t}\Big)dt$
$=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{|\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}-e^{-t}\Big)dt,$
where the first equation is from (9.5). Therefore, we have
$\log\zeta(0)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r)=\log\rho_{{\rm R}_{0}f}(u),$
which concludes the proof. ∎
9.2. Results for $s$-concave functions
In this subsection, we focus on the case of $s$-concave functions. The following lemma is a direct corollary from the Borell-Brascamp-Lieb inequality, which is also mentioned in [15, Theorem 11.2].
Lemma 9.7.
For $s$-concave $f\in L^{1}(\mathbb{R}^{n})$ and fixed $u\in\mathbb{S}^{n-1}$,
$t\mapsto g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu))\}dx$
is a $s/(ns+1)$-concave function on $\mathbb{R}$.
Proof.
We denote by
$N_{f}(x,t)=\min\{f(x),f(x+tu)\}.$
First, we prove that $N_{f}(x,t)$ is $s$-concave on $\mathbb{R}^{n}\times\mathbb{R}$. For $x_{0},x_{1}\in\mathbb{R}^{n}$, $t_{0},t_{1}\in\mathbb{R}$ and $\lambda\in[0,1]$, we denote $x_{\lambda}=(1-\lambda)x_{0}+\lambda x_{1}$ and $t_{\lambda}=(1-\lambda)t_{0}+\lambda t_{1}$. Note that
$N_{f}(x_{\lambda},t_{\lambda})^{s}=\min\{f(x_{\lambda})^{s},f(x_{\lambda}+t_{\lambda}u)^{s}\}.$
Since $f$ is $s$-concave and $\min\{\cdot,\cdot\}$ is a concave function on $\mathbb{R}\times\mathbb{R}$, we have
$N_{f}(x_{\lambda},t_{\lambda})^{s}$ $\geq\min\{(1-\lambda)f(x_{0})^{s}+\lambda f(x_{1})^{s},(1-\lambda)f(x_{0}+t_{0}u)^{s}+\lambda f(x_{1}+t_{1}u)^{s}\}$
$\geq(1-\lambda)\min\{f(x_{0})^{s},f(x_{0}+t_{0}u)^{s}\}+\lambda\min\{f(x_{1})^{s},f(x_{1}+t_{1}u)^{s}\}$
$=(1-\lambda)N_{f}(x_{0},t_{0})^{s}+\lambda N_{f}(x_{1},t_{1})^{s},$
which implies the $s$-concavity of $N_{f}$.
For $\lambda\in[0,1]$, and fixed $t_{0},t_{1}\in\mathbb{R}$, we set
$F(x)=N_{f}(x,t_{0}),~~G(x)=N_{f}(x,t_{1}),~~H(x)=N_{f}(x,t_{\lambda}),$
where $t_{\lambda}=(1-\lambda)t_{0}+\lambda t_{1}$. By the $s$-concavity of $F$, we have
$H((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(F(x),G(y)).$
The Borell-Brascamp-Lieb inequality (2.3) then implies that
$\int_{\mathbb{R}^{n}}H\geq\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}F,\int_{\mathbb{R}^{n}}G\Big),$ (9.6)
that is, the $\frac{s}{ns+1}$-concavity of $g(tu)$ in $t$. ∎
We begin the proof of Theorem 5 by stating its inclusion result as the following theorem; the equality case will be addressed separately.
Theorem 9.8.
Let $f\in L^{1}(\mathbb{R}^{n})$ be $s$-concave with compact support. If $1<\beta<\alpha$, then
${\rm R}_{\infty}f\subset c_{n,\beta}(s){\rm R}_{\beta}f\subset c_{n,\alpha}(s){\rm R}_{\alpha}f\subset(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$ (9.7)
where $c_{n,\alpha}(s)=((n+\frac{1}{s})B(\alpha+1,n+\frac{1}{s}))^{-1/\alpha}$.
Proof.
By Lemma 9.7, the function $g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ is $s^{\prime}$-concave with $s^{\prime}=s/(ns+1)$. We apply Lemma 9.2 with $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$.
Therefore, we obtain that
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}g(tu)dt}{\int_{0}^{\infty}g(0)t^{\alpha-1}(1-s^{\prime}t)_{+}^{1/s^{\prime}}dt}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
for $\alpha>0$ and
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}(g(0)-g(tu))dt}{\int_{0}^{\infty}g(0)t^{\alpha-1}((1-s^{\prime}t)_{+}^{1/s^{\prime}}-1)dt}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
for $1<\alpha<0$, where we use (9.1) and (9.3). Note that
$\Big(\alpha B(\alpha,1+\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=\Big(nB(\alpha+1,\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=c_{n,\alpha}(s).$
Lemma 9.2 then implies that
$c_{n,\beta}(s){\rm R}_{\beta}f\subset c_{n,\alpha}(s){\rm R}_{\alpha}f,$ (9.8)
for $1<\alpha<\beta$.
Since Lemma 8.3 ensures that ${\rm R}_{\beta}f\to{\rm R}_{\infty}f$ and $c_{n,\beta}(s)\to 1$ as $\beta\rightarrow\infty$, the left inclusion follows directly from (9.8).
By Lemma 8.2 and $c_{n,\alpha}(s)\sim\frac{\alpha+1}{n+1/s}$ when $\alpha\rightarrow-1$, we have
$\lim_{\alpha\rightarrow-1^{+}}c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u)=(n+\frac{1}{s})\|f\|_{1}\rho_{\Pi^{*}f}(u).$
which yields the right inclusion. ∎
When $s=\infty$, Theorem 9.8 reduces to the geometric case in Theorem 9.1 by Gardner and Zhang. In fact, following the argument of our proof, for a log-concave function $f\in L^{1}(\mathbb{R}^{n})$ with compact support, if one could show that $g(tu)^{1/n}$ is concave, then the following chain of inclusions would hold:
${\rm R}_{\infty}f\subset c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f\subset n\|f\|_{1}\Pi^{*}f,$
for $1<\alpha<\beta<\infty$.
This would yield precisely the functional counterpart of Theorem 9.1 that one would expect. However, $g(tu)^{1/n}$ is not necessarily concave for a general log-concave function $f$. A counterexample will be provided below.
Let $n=1$ and
[…]
It is clear that $\log f(x)=\log x+\log 2$ for $t\in[0,1]$, which is concave. Then we have
[…]
Note that
$\frac{1}{4}=g(\frac{1}{2})<\frac{1}{2}(g(0)+g(1))=\frac{1}{2},$
which shows that $g$ is not concave. In this case, we consider $\beta=1$ and $\alpha=1/2$ and obtain that
$c_{1,1}{\rm R}_{1}f=\Big[-\frac{2}{3},\frac{2}{3}\Big],\quad c_{1,\frac{1}{2}}{\rm R}_{\frac{1}{2}}f=\Big[-\frac{16}{25},\frac{16}{{25}}\Big].$
which is not consistent with the desired result $c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f$.
The following theorem describes the sharpness of the inclusion result in Theorem 5.
Theorem 9.9.
Equality holds in (9.7) if $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$ where $x_{0}\in\mathbb{R}^{n}$, $c>0$ and $\triangle$ is a simplex containing the origin.
Proof.
Note that ${\rm R}_{\alpha}f$ is invariant under scaling and translation of $f$. Without loss of generality, we may assume $f(x)=(1-\|x\|_{\triangle})_{+}^{1/s}$. A direct computation then gives
$\|f\|_{1}=B(\frac{1}{s},n+1)\frac{|\triangle|}{s}.$
We use the same notation for $g(tu)$ with a fixed $u\in\mathbb{S}^{n-1}$, and we have
$g(tu)$ $=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
$=\int_{0}^{1}|\{f\geq r\}\cap(\{f\geq r\}+tu)|dr$
$=\int_{0}^{1}\frac{1}{s}r^{\frac{1}{s}-1}|(1-r)\triangle\cap((1-r)\triangle+tu)|dr$
$=\int_{0}^{1}\frac{1}{s}r^{\frac{1}{s}-1}(1-r)^{n}|\triangle\cap(\triangle+\frac{t}{1-r}u)|dr.$
By [18, Lemma 3.4], $|\triangle\cap(\triangle+tu)|^{1/n}$ is linear in $t$. Therefore, there exists $\lambda>0$ such that
$|\triangle\cap(\triangle+\frac{t}{1-r})u|=|\triangle|(1-\lambda\frac{t}{1-r})_{+}^{n}.$
A straightforward, though somewhat tedious, calculation then shows that
$g(tu)=\frac{1}{s}B(\frac{1}{s},n+1)|\triangle|(1-\lambda t)^{\frac{ns+1}{s}}.$
We set $s^{\prime}=s/(ns+1)$. Then, for $1<\alpha<0$ and $\alpha>0$, by taking $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$ in Lemma 9.2, we obtain
$c_{n,\alpha}(s){\rm R}_{\alpha}f=\frac{1}{s^{\prime}}\zeta(\alpha)=\frac{1}{s^{\prime}}\Big(\frac{\int_{0}^{\infty}t^{\alpha-1}(1-\lambda t)_{+}^{1/s^{\prime}}dt}{\int_{0}^{\infty}t^{\alpha-1}(1-s^{\prime}t)^{1/s^{\prime}}dt}\Big)^{1/\alpha}=\frac{1}{\lambda},$
where the first equality is from thex calculation in Theorem 9.8.
This implies that
$c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f,\quad\quad-1<\alpha<\beta<\infty.$
Letting $\alpha\rightarrow-1$ and $\beta\rightarrow\infty$, we have
${\rm R}_{\infty}f=c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f=(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$
which concludes the proof. ∎
We conclude the section by showing the necessary condition for the equality cases in Theorem 5. The following lemma is the key result for the equality characterization.
Lemma 9.10.
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ non-zero and $s$-concave with compact support. If
$g(tu)^{\frac{s}{ns+1}}=\Big(\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx\Big)^{\frac{s}{ns+1}}$
is linear in $t$ for any fixed $u\in\mathbb{S}^{n-1}$, then $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$, where $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.
Proof.
We assume, without loss of generality, that $\|f\|_{\infty}=1$. Since
$g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x-tu)\}dx,$
for convenience of notations in the following proof, we set $N_{f}(x,t)=\min\{f(x),f(x-tu)\}$ here. The functions $F,G,H$ are as defined by
$F(x)=N_{f}(x,0),~~G(x)=N_{f}(x,t),~~H(x)=N_{f}(x,\lambda t),$
which satisfies that $H((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(F(x),G(y))$.
The linearity of $g(tu)^{\frac{s}{ns+1}}$ then implies that
$\int_{\mathbb{R}^{n}}H=\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}F,\int_{\mathbb{R}^{n}}G\Big).$
That is, we have the equality in the Borell-Brascamp-Lieb inequality for $F,G,H$.
By Theorem 2.4, there exists $c_{0}=\Big(\frac{|{\rm supp}(G)|}{|{\rm supp}(F)|}\Big)^{1/n}$ and $y_{0}\in\mathbb{R}^{n}$ such that
${\rm supp}(G)=c_{0}\,{\rm supp}(F)+y_{0},$ (9.9)
while the definition of $G$ shows
${\rm supp}(G)={\rm supp}(F)\cap({\rm supp}(F)+tu).$ (9.10)
Since $u\in\mathbb{S}^{n-1}$ is arbitrary fixed, (9.9) and (9.10) shows that ${\rm supp}(F)\cap({\rm supp}(F)+tu)$ is homothetic to ${\rm supp}(F)$ for all $u\in\mathbb{S}^{n-1}$. By [18, Lemma 3.4], we have ${\rm supp}(F)$ is a simplex $\triangle$, where we assume that $|\triangle|$ is the same as the volume of the standard simplex $\triangle_{n}$ for convenience.
In fact, we only need to consider when $\triangle$ is the standard simplex, that is,
$\triangle=\triangle_{n}=\Big\{x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}:x_{i}\geq 0,~~\sum_{i=1}^{n}x_{i}\leq 1\Big\}.$
Otherwise, there exists $\phi\in{\rm SL}(n)$ such that ${\rm supp}(f\circ\phi^{-1})=\triangle_{n}$ and we consider $f\circ\phi^{-1}$ instead of $f$. In this case, by (9.9) and (9.10), we have
$c_{0}(t)=1-t\sum_{i=1}^{n}u_{i},~~\text{and}~~y_{0}(t)=t(u_{1}^{+},\ldots,u_{n}^{+}),$
where $u_{i}^{+}=\max\{0,u_{i}\}$ and we denote by $a_{0}=\sum_{i=1}^{n}u_{i}$. Therefore, we can choose $u\in\mathbb{S}^{n-1}$ with $u_{i}\geq 0$, such that $y_{0}(t)=tu$.
Moreover, by Theorem 2.4, there exists a $(\lambda_{0},s)$-concave function $\Phi:{\rm supp}(F)\rightarrow[0,\infty)$ with $\lambda_{0}=\frac{\lambda c_{0}}{1-\lambda+\lambda c_{0}}$ satisfying,
$\left\{\begin{aligned} &F(x)=\Phi(x)&&,\\ &G(c_{0}x+tu)=c_{0}^{\frac{1}{s}}\Phi(x)&&,\\ &H((1-\lambda+\lambda c_{0})x+\lambda tu)=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{0}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&.\end{aligned}\right$.
In fact, $\Phi(x)=F(x)=f(x)$ by the definition of $F$. Then we have
$G((1-a_{0}t)x+tu)^{s}=\min\{f((1-a_{0}t)x)^{s},f((1-a_{0}t)x+tu)^{s}\}=(1-a_{0}t)f(x)^{s}.$
Note that $(1-a_{0}t)f(x)^{s}\neq f((1-a_{0}t)x)^{s}$ since $f^{s}$ is concave. Therefore,
$f((1-a_{o}t)x+tu)^{s}=(1-a_{0}t)f(x)^{s}$ (9.11)
holds for $0\leq t\leq 1/\sum u_{i}$. Equation (9.11), together with $f\geq 0$ and the $s$-concavity of $f$ , imply that
$f\Big(\frac{u}{a_{0}}\Big)^{s}=0.$ (9.12)
We set $\bar{x}=(\bar{x}_{1},\ldots,\bar{x}_{n})\in\triangle_{n}$ such that $f(\bar{x})=\|f\|_{\infty}=1$. Note that if $\sum_{i}\bar{x}_{i}\in(0,1)$, we will have
$f\Big(\frac{\bar{x}}{\sum_{i}\bar{x}_{i}}\Big)^{s}=0\geq(1-\sum_{i}\bar{x}_{i})f(\bar{x})^{s}+\Big(\sum_{i}\bar{x}_{i}\Big)f(o)^{s}\geq 1-\sum_{i}\bar{x}_{i}>0,$ (9.13)
which cause a contradiction. Thus we have $\sum_{i}\bar{x}_{i}=1$ or $\bar{x}=o$. If $\sum_{i}\bar{x}_{i}=1$, equation (9.12) implies that $f(\bar{x})=0$ which contradicts to the choice of $\bar{x}$. Thus we have $\bar{x}=o$. Letting $x=\bar{x}=o$ in (9.11) then gives
$f(tu)^{s}=1-a_{0}t=1-\|tu\|_{\triangle_{n}}.$
Since ${\rm supp}(f)=\triangle_{n}$, we then have
$f(x)=(1-\|x\|_{\triangle_{n}})_{+}^{1/s},$
which concludes the proof. ∎
We conclude this section by the following theorem, which provides the full equality characterization in Theorem 5.
Theorem 9.11.
Let $s>0$ and $c_{n,\alpha}(s)$ be defined in Theorem 5. Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, and $s$-concave function with compact support. For $1<\alpha<\beta<\infty$, if
$c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f,$
then $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$ where $c>0$, $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.
Proof.
Let $g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ with fixed $u\in\mathbb{S}^{n-1}$. By Theorem 9.8, letting $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$ in Lemma 9.2, we obtain
$\zeta(\alpha)=s^{\prime}c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u),$
where $s^{\prime}=s/(ns+1)$. Then $c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f$ implies that $\zeta(\alpha)=\zeta(\beta)$. Since $\omega$ is strictly decreasing on its support, by Lemma 9.3, we obtain $\lambda=\zeta(\alpha)^{-1}$ such that
$\varphi(t)=\frac{1}{s^{\prime}}\Big(1-\Big(\frac{g(tu)}{g(0)}\Big)^{s^{\prime}}\Big)=\lambda t.$
This implies that $g(tu)^{s^{\prime}}$ is linear in $t$. Then we conclude the proof by Lemma 9.10. ∎
Acknowledgements: This research was funded in whole or in part by the Austrian Science Fund (FWF) doi/10.55776/37030. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.