### v2

# Affine chord Sobolev inequalities and radial mean bodies for functions

Abstract.
Affine isoperimetric inequalities for the functional radial mean bodies are derived from the new affine chord Sobolev inequalities, which extend the recent affine isoperimetric inequalities of Haddad and Ludwig [24] from convex bodies to functions. The affine chord Sobolev inequalities further imply a strengthening of the Euclidean chord Sobolev inequalities introduced in [4]. Moreover, for $s$-concave functions $f$ with compact support and $s>0$, a parameter-dependent monotonicity property of the functional radial mean body ${\rm R}_{\alpha}f$ is obtained: ${\rm R}_{\alpha}f\subset{\rm R}_{\beta}f$ for $1<\alpha<\beta$, and, after suitable normalization, the reverse inclusion also holds. These sharp results generalize the corresponding monotonicity for geometric radial mean bodies established by Gardner and Zhang [18].
1. Introduction
Affine inequalities, invariant under translations and volume-preserving linear transformations, often strengthen classical Euclidean inequalities. A seminal example is Zhang’s affine Sobolev inequality [48], which significantly improves the classical Sobolev inequality. Wang [44] later extended this inequality to functions of bounded variation, yielding a generalized Petty projection inequality (see also [48]), itself a strengthening of the classical isoperimetric inequality. Recently, Milman and Yehudayoff [39] established affine isoperimetric inequalities for affine quermassintegrals, settling a long-standing conjecture of Lutwak [33]. For further developments on affine inequalities, see, for example, [7, 10, 12, 18, 20, 21, 22, 23, 34, 35, 36].
More recently, Haddad and Ludwig [24, 25, 26] developed a new approach to study affine inequalities. They proved the following sharp affine fractional Sobolev inequalities in [24], which are stronger than their Euclidean counterparts by Almgren and Lieb [2] (see also [15]).
Affine fractional Sobolev inequalities: For $\alpha\in(-1,0)$ and $f\in W^{-\alpha,1}(\mathbb{R}^{n})$,
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}$ $r^{\alpha-1}\|f(\cdot)-f(\cdot+ru)\|_{1}dr\Big)^{\frac{n}{\alpha}}du\Big)^{\frac{\alpha}{n}}$ (1.1)
$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,$
where $\mathbb{S}^{n-1}$ is the unit sphere in $\mathbb{R}^{n}$ and $\omega_{n}$ denotes the volume of the unit ball in $\mathbb{R}^{n}$. The sharp constant $\sigma_{n,\alpha}$ is given by
$\sigma_{n,\alpha}=\frac{n2^{\alpha}\pi^{\frac{n-\alpha-1}{2}}\Gamma(\frac{\alpha+1}{2})\Gamma(\frac{n}{2}+1)^{\frac{\alpha}{n}}}{|\alpha|\Gamma(\frac{n+\alpha+1}{2})}.$ (1.2)
Here $\Gamma$ denotes the gamma function, $\|f\|_{p}:=(\int_{\mathbb{R}^{n}}f(x)^{p}dx)^{1/p}$ denotes the $L^{p}$ norm of $f$ and $W^{-\alpha,1}(\mathbb{R}^{n})$ is the space of integrable functions $f$ such that the right-most side of (1.1) is finite.
Eliminating the intermediate term in (1.1) gives the fractional Sobolev inequalities, which converge to the classical Sobolev inequality when $\alpha\rightarrow-1^{+}$, as shown by Bourgain, Brezis and Mironescu [8]. Haddad and Ludwig [24] showed that (1.1) becomes the affine Sobolev inequality in the limit $\alpha\rightarrow-1^{+}$. The fractional Sobolev inequalities have been extensively studied over the past decades (see, e.g., [1, 8, 13, 27, 30, 31]). In very recent work by the authors [4], the following sharp chord Sobolev inequalities for positive $\alpha$ are established, where, for $E\subset\mathbb{R}^{n}$, the characteristic function $\operatorname{1}_{E}(x)$ takes value $1$ with $x\in E$ and 0 otherwise.
Chord Sobolev inequalities for $\boldsymbol{\alpha>0}$: For $\alpha\in(0,n)$ and non-negative $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$ (1.3)
For $\alpha>n$ and non-zero, non-negative $f\in L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy,$ (1.4)
where $\sigma_{n,\alpha}$ is given by (1.2). In both cases, equality holds if and only if $f$ is a constant multiple of the characteristic function of a ball.
In [4], fractional Sobolev inequalities, together with (1.3) and (1.4), form a complete series called the chord Sobolev inequalities. When $f=\operatorname{1}_{K}$ is the characteristic function of a convex body $K\subset\mathbb{R}^{n}$ (a convex compact set with non-empty interior), these inequalities reduce to the isoperimetric inequalities for chord power integrals; see [4] for details. Chord power integrals have been studied extensively in recent years [4, 11, 19, 37, 45, 46].
One of the main aims of this paper is to prove the following affine chord Sobolev inequalities for $\alpha>0$, which strengthen their Euclidean counterparts (1.3) and (1.4). Here $f$ is radially symmetric if its value depends only on $|x|$.
Theorem 1.
For $\alpha\in(0,n)$ and non-negative $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}$ $\big\|\min\{f(\cdot),f(\cdot+ru)\}\big\|_{1}dr\Big)^{\frac{n}{\alpha}}du\Big)^{\frac{\alpha}{n}}$ (1.5)
$\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
There is equality in the first inequality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid. There is equality in the second inequality if $f$ is radially symmetric.
Theorem 2.
For $\alpha>n$ and non-zero, non-negative $f\in L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}$ $\big\|\min\{f(\cdot),f(\cdot+ru)\}\big\|_{1}dr\Big)^{\frac{n}{\alpha}}du\Big)^{\frac{\alpha}{n}}$ (1.6)
$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
There is equality in the first inequality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid. There is equality in the second inequality if $f$ is radially symmetric.
The affine fractional Sobolev inequalities (1.1), together with (1.5) and (1.6), are called affine chord Sobolev inequalities. To prove Theorem 1 and Theorem 2, we introduce an anisotropic version of the right-most side of (1.5) (see, e.g., [24, 30, 31]), which gives rise to a new star-shaped set ${\rm L}_{\alpha}f$, defined via its radial function by
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr,~~u\in\mathbb{S}^{n-1},$
(see Section 3 for details). The first inequality in Theorem 1 can then be written as
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}},$ (1.7)
where $|{\rm L}_{\alpha}f|$ denotes the volume of ${\rm L}_{\alpha}f$. Since both sides of (1.7) are invariant under translations of $f$ and
${\rm L}_{\alpha}(f\circ\phi^{-1})=\phi{\rm L}_{\alpha}f$
for every volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, it follows that inequality (1.7) is affine invariant.
For a convex body $K\subset\mathbb{R}^{n}$, the star-shaped set ${\rm L}_{\alpha}\operatorname{1}_{K}$ is proportional to the radial mean body ${\rm R}_{\alpha}K$ of Gardner and Zhang [18], as will be discussed in Section 5. Haddad and Ludwig [24] proved affine isoperimetric inequalities for ${\rm R}_{\alpha}K$ with $\alpha>-1$, stating that
$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha\in(-1,n),$ (1.8)
$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha>n.$
Equality holds if and only if $K$ is an ellipsoid.
In [4], the authors extended the notion ${\rm R}_{\alpha}K$ to log-concave functions $f\in L^{1}(\mathbb{R}^{n})$ by
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t),~\alpha\neq 0,$ (1.9)
and
$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq t\}}(u)d\mu_{f}(t),$
where $d\mu_{f}(t)=\frac{|\{f\geq t\}|}{\|f\|_{1}}dt$ is a probability measure on $\mathbb{R}$. We note that Haddad and Ludwig [26] also use the notation ${\rm R}_{\alpha}f$ for the $\alpha$-th radial mean body, but with a different definition. The present definition can be viewed as an $L^{1}$ extension to their definitions.
It was shown in [4] that ${\rm R}_{\alpha}f$ is proportional to ${\rm L}_{\alpha}f$ for $\alpha>0$, while for $\alpha\in(-1,0)$ it is proportional to the fractional polar projection body $\Pi^{*,-\alpha}f$ arising from the affine fractional Sobolev inequalities (1.1). Consequently, the affine chord Sobolev inequalities yield the following affine isoperimetric inequalities for ${\rm R}_{\alpha}f$: for a log-concave function $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$,
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha\in(-1,n),$ (1.10)
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{1}\|f\|_{\infty}^{-1}}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha>n,$
with equality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid. In particular, when $f=\operatorname{1}_{K}$ is the characteristic function of a convex body, (1.10) reduces to the geometric inequality (1.8).
Of particular interest, for log-concave $f\in L^{1}(\mathbb{R}^{n})$, Haddad and Ludwig [24] proved
$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|f\|_{1}\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\rho_{\Pi^{*}f}(u)^{-1},$
where $\Pi^{*}f$ is the polar projection body of $f$ introduced by Zhang [48] and extended by Wang [44] (see Section 2.5). For the other endpoint $\alpha\to\infty$, the monotonicity of $L^{p}$ norms with respect to $p$, together with (1.9), leads to the definition of ${\rm R}_{\infty}f$:
${\rm R}_{\infty}f={\rm R}_{\infty}\,{\rm supp}(f)={\rm D}\,{\rm supp}(f),$
where ${\rm D}K:=\{x-y:x,y\in K\}$ denotes the difference body, and we set ${\rm R}_{\infty}f=\mathbb{R}^{n}$ if ${\rm supp}(f)=\mathbb{R}^{n}$.
In Section 5, we prove the following monotonicity for log-concave $f\in L^{1}(\mathbb{R}^{n})$:
${\rm R}_{\alpha}f\subset{\rm R}_{\beta}f\subset{\rm R}_{\infty}f$
and
$\|f\|_{\infty}\Pi^{*}f\subset\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$ (1.11)
for $1<\alpha<\beta<\infty$. When $f=\operatorname{1}_{K}$, these reduce to the geometric case of Gardner and Zhang [18].
Gardner and Zhang [18] also established the reverse inclusion for radial mean bodies, which is more involved. For a convex body $K\subset\mathbb{R}^{n}$, they proved
${\rm D}K\subset c_{n,\beta}{\rm R}_{\beta}K\subset c_{n,\alpha}{\rm R}_{\alpha}K\subset n|K|\Pi^{*}K,$ (1.12)
for $1<\alpha<\beta<\infty$, where $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$ and $B(\cdot,\cdot)$ is the Beta function. Equality holds in each inclusion if and only if $K$ is a simplex. When $\beta=n$, the left-most inclusion gives the Rogers–Shephard inequality [41], while for $\alpha=n$, the right-most inclusion yields Zhang’s inequality [47] (see also Ball [5]).
Building on the functional radial mean body ${\rm R}_{\alpha}f$ and (1.11), we aim to extend the reverse inclusion (1.12) from sets to functions. There are two natural ways to associate a convex body with a log-concave function:
$K\mapsto\operatorname{1}_{K}\quad\quad\text{and}\quad\quad K\mapsto e^{-\|\cdot\|_{K}}.$
Using the second correspondence, Langharst, Marín Sola and Ulivelli [28] showed that for log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f\subset 2\|f\|_{1}\Pi^{*}f.$ (1.13)
Equality holds in each inclusion if and only if $f(x)=\|f\|_{\infty}e^{-\|x-x_{0}\|_{\triangle}}$, where $x_{0}$ is a maximizer of $f$ and $\triangle$ is a simplex containing the origin and $x_{0}$. When $\alpha=n$, the right-most inclusion recovers Zhang’s inequality for log-concave functions, as established by Alonso-Gutiérrez, Bernués and González Merino [3]. However, in the limit $\beta\rightarrow\infty$, the body ${\rm R}_{\infty}f$ no longer appears in (1.13). This indicates a limitation of the second correspondence in capturing the endpoint behavior of radial mean bodies. On the other hand, the monotonicity property (1.11) naturally extends from convex bodies to functions via the embedding $K\mapsto\operatorname{1}_{K}$. These considerations motivate us to adopt the first correspondence and lead to the following question.
Question. For a non-zero, log-concave function $f\in L^{1}(\mathbb{R}^{n})$ with compact support, does
${\rm R}_{\infty}f\subset c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f\subset n\|f\|_{1}\Pi^{*}f$
hold for $1<\alpha<\beta<\infty$? Here $c_{n,\alpha}$ is the constant in (1.12).
Unfortunately, this inclusion fails for general log-concave functions, as we show by a counterexample in Section 6. This failure leads us to consider $s$-concave functions with $s>0$.
Let $f_{+}(x):=\max\{f(x),0\}$ denote its positive part. The following theorem extends (1.12) to $s$-concave functions, with fully characterized equality cases.
Theorem 3.
Let $s>0$ and let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. For $1<\alpha<\beta<\infty$, the following chain of inclusions holds:
${\rm R}_{\infty}f\subset c_{n,\beta}(s)\,{\rm R}_{\beta}f\subset c_{n,\alpha}(s)\,{\rm R}_{\alpha}f\subset\Big(n+\frac{1}{s}\Big)\|f\|_{1}\,\Pi^{*}f,$ (1.14)
where $c_{n,\alpha}(s)=\bigl((n+\tfrac{1}{s})\,B(\alpha+1,n+\tfrac{1}{s})\bigr)^{-1/\alpha}$. In each inclusion, there is equality if and only if $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
Letting $\beta=n$ and $\alpha=n$ in (1.14) yields the Rogers-Shephard inequality and Zhang inequality, respectively, for $s$-concave functions; see Section 6 for detailed statement.
The optimizer in (1.14) admits a simple geometric interpretation. Let $\triangle$ be a simplex in $\mathbb{R}^{n}$ with vertices $v_{0},\ldots,v_{n}$, and identify $\mathbb{R}^{n}$ with the hyperplane $\mathbb{R}^{n}\times\{0\}\subset\mathbb{R}^{n+1}$. For $f^{s}(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}\quad\text{with }x_{0}\in\triangle$, its hypograph is given by the set
$\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm supp}(f^{s}),~0\leq l\leq f^{s}(x)\}$
and coincides exactly with the simplex in $\mathbb{R}^{n+1}$ having vertices $v_{0},\ldots,v_{n},(x_{0},a)$. The equality characterization of (1.14) relies on the equality case of the Borell–Brascamp–Lieb inequality. Its highly rigid structure accounts for the difficulty in the study of the equality cases.
2. Preliminaries
In this section, we collect basic notation and results on function spaces and Schwarz symmetrals, together with some basic geometric facts. Useful general references include the books by Gardner [16], Schneider [42], and Schneider and Weil [43].
2.1. $L^{p}$ spaces
Let p>1p>-1 and $(X,\mathcal{B},\mu)$ be a measure space. In this work, we primarily focus on the cases $X=\mathbb{R}^{n}$ or $X\subset\mathbb{R}^{n}$ with $\mu$ being the $n$-dimensional Lebesgue measure.
For $p\neq 0$ and a measurable function $f:X\rightarrow\mathbb{R}$, let
$\|f\|_{p}=\Big(\int_{X}|f(x)|^{p}d\mu(x)\Big)^{1/p},$
and
$\|f\|_{\infty}=\inf\{C>0:|f(x)|\leq C~~\text{almost everywhere on}~X\}.$
In particular, for $p\in(-1,0)$, we set $\|f\|_{p}=\infty$ if $\mu(\{x\in X:f(x)=0\})>0$. The associated space $L^{p}(X)$ is then defined as
$L^{p}(X)=\{f:X\rightarrow\mathbb{R}:f~\text{is measurable},~\|f\|_{p}<\infty\},$
where $f=g$ in $L^{p}(X)$ means that they are equal almost everywhere.
When $\mu(X)=1$, the $L^{p}$ norm is monotone in $p$: for 1<p<q<0-1<p<q<0 or $0<p<q<\infty$,
$\|f\|_{p}\leq\|f\|_{q},$
as a consequence of Jensen’s inequality. In fact, this monotonicity also extends for $p<0<q$. To see this, under the assumption $\mu(X)=1$, we first note the limit
$\lim_{p\rightarrow 0}\log\|f\|_{p}=\int_{X}\log|f(x)|d\mu(x),$
which follows from the continuity of $\|f\|_{p}$ with respect to $p$. Applying Jensen’s inequality then gives
$\|f\|_{p}\leq\int_{X}\log|f(x)|d\mu(x)\leq\|f\|_{q},$
for $p<0<q$.
2.2. Symmetrization
For a Borel set $E\subset\mathbb{R}^{n}$ with finite measure, the Schwarz symmetral of $E$, denoted by $E^{\star}$, is the closed, centered ball having the same volume as $E$. Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function. For $t>0$, the superlevel set of $f$ is defined by $\{f\geq t\}=\{x\in\mathbb{R}^{n}:f(x)\geq t\}$. We say $f$ is non-zero if $\{f\neq 0\}$ has positive measure.
If the superlevel sets of $f$ have finite measure, the layer cake formula gives
$f(x)=\int_{0}^{\infty}\operatorname{1}_{\{f\geq t\}}(x)dt$
for almost every $x\in\mathbb{R}^{n}$. The Schwarz symmetral of $f$, denoted by $f^{\star}$, is defined as
$f^{\star}(x)=\int_{0}^{\infty}\operatorname{1}_{\{f\geq t\}^{\star}}(x)dt$
for $x\in\mathbb{R}^{n}$, so that $f^{\star}$ is radially symmetric and the superlevel sets satisfy
$|\{f^{\star}\geq t\}|=|\{f\geq t\}|$
for all $t>0$.
A central tool in our proofs is the Riesz rearrangement inequality (see, for example [29, Theorem 3.7]) together with the characterization of equality by Burchard [9].
Theorem 2.1 (Riesz’s rearrangement inequality).
For measurable $f,g,k:\mathbb{R}^{n}\rightarrow[0,\infty)$ with superlevel sets of finite measure.
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}f(x)k(x-y)g(y)dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}f^{\star}(x)k^{\star}(x-y)g^{\star}(y)dxdy.$
Theorem 2.2 (Burchard).
Let $A,B$ and $C$ be sets of finite positive measure in $\mathbb{R}^{n}$ and denote by $a,b$ and $c$ the radii of their Schwarz symmetrals $A^{\star},B^{\star}$ and $C^{\star}$. For $|\alpha-\beta|<\gamma<\alpha+\beta$, there is equality in
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{A}(x)\operatorname{1}_{B}(x-y)\operatorname{1}_{C}(y)dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{A^{\star}}(x)\operatorname{1}_{B^{\star}}(x-y)\operatorname{1}_{C^{\star}}(y)dxdy$
if and only if, up to sets of measure zero,
$A=x_{0}+aD,~B=x_{1}+bD,~C=x_{2}+cD,$
where $D$ is a centered ellipsoid, and $x_{0},x_{1}$ and $x_{2}=x_{0}+x_{1}$ are vectors in $\mathbb{R}^{n}$.
2.3. Star-shaped sets and dual mixed volumes
A set $K\subset\mathbb{R}^{n}$ is called star-shaped (with respect to the origin) if for every $x\in K$, the line segment $[o,x]$ is contained in $K$. For a star-shaped set $K$ containing the origin in its interior, the gauge function $\|\cdot\|_{K}:\mathbb{R}^{n}\rightarrow[0,\infty]$ is defined by
$\|x\|_{K}=\inf\{\lambda>0:x\in\lambda K\}.$ (2.1)
If $K$ contains the origin, possibly on its boundary, then for any point $x$, the gauge function $\|x\|_{K}$ is defined by (2.1) if there exists $r>0$ such that $rx\in K$; otherwise, $\|x\|_{K}=\infty$.
[…]
$\rho_{K}(x)=\|x\|_{K}^{-1}=\sup\{r\geq 0:rx\in K\}.$
In particular, if $K$ is the unit Euclidean ball, $\|\cdot\|_{K}$ coincides with the Euclidean norm $|\cdot|$.
For a star-shaped set $K$ with measurable radial function, the volume, or the $n$-dimensional Lebesgue measure of $K$ is given by
$|K|=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n}du,$
where $du$ denotes the spherical Lebesgue measure. Following Lutwak [32], the dual mixed volume for star-shaped sets $K$ and $L$ is defined by
$\tilde{V}_{\alpha}(K,L)=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\rho_{L}(u)^{\alpha}du,$
[…]
$\tilde{V}_{\log}(K,L)=\lim_{\alpha\rightarrow 0}\log\Big(\frac{\tilde{V}_{\alpha}(K,L)}{|K|}\Big)^{1/\alpha}=\frac{1}{n|K|}\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n}\log\Big(\frac{\rho_{K}(u)}{\rho_{L}(u)}\Big)du.$
If $\alpha\in(0,n)$ and $K,L$ are star-shaped sets with finite volume, the dual mixed volume inequality states that
$\tilde{V}_{\alpha}(K,L)\leq|K|^{\frac{n-\alpha}{n}}|L|^{\frac{\alpha}{n}},$ (2.2)
with equality holds if and only if $K$ and $L$ are dilates, i.e., $\rho_{K}(u)=c\rho_{L}(u)$ for some $c\geq 0$ and almost all $u\in\mathbb{S}^{n-1}$. For $\alpha<0$ or $\alpha>n$, the dual mixed volume inequality states that
$\tilde{V}_{\alpha}(K,L)\geq|K|^{\frac{n-\alpha}{n}}|L|^{\frac{\alpha}{n}},$ (2.3)
where equality holds if and only if $K$ and $L$ are dilates.
2.4. Log-concave and $s$-concave functions
A function $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is log-concave if $\log f$ is a concave function on $\mathbb{R}^{n}$, with values in $[-\infty,\infty)$. For $s>0$, a function $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is $s$-concave if $f^{s}$ is concave on its support.
Lemma 2.3.
[3, Lemma 2.1] Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a log-concave and integrable function. Then the function $g:\mathbb{R}^{n}\rightarrow[0,\infty)$ defined by
$g(x)=\int_{\mathbb{R}^{n}}\min\{f(y),f(y+x)\}dy$
is even, log-concave, with $\|g\|_{\infty}=g(o)=\|f\|_{1}$ and ${\rm supp}(g)$ containing the origin point in its interior.
For $\lambda\in[0,1]$, $a,b>0$ and $s>0$, we denote by
[…]
For an $s$-concave function $f$, we have $f((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(f(x),f(y))$ for $\lambda\in[0,1]$. We say $f$ is $(\lambda_{0},s)$-concave, if $f((1-\lambda_{0})x+\lambda_{0}y)\geq\mathcal{M}_{\lambda_{0}}^{s}(f(x),f(y))$ for fixed $\lambda_{0}\in(0,1)$.
The following theorem of Balogh and Kristály [6] gives the equality characterization of the Borell-Brascamp-Lieb inequality.
Theorem 2.4.
[6, Theorem 3.1] Let $\lambda\in(0,1)$ and $s>-\frac{1}{n}$. Let $F,G,H:\mathbb{R}^{n}\rightarrow[0,\infty)$ be non-zero, compactly supported integrable functions satisfying
$H((1-\lambda)x+\lambda y)\geq{\mathcal{M}}_{\lambda}^{s}(F(x),G(y)).$
Then there is equality in
$\int_{\mathbb{R}^{n}}H\geq{\mathcal{M}}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}F,\int_{\mathbb{R}^{n}}G\Big),$ (2.4)
if and only if there is $x_{0}\in\mathbb{R}^{n}$, an $(\lambda_{0},s)$-concave function $\Phi:{\rm supp}(F)\rightarrow[0,\infty)$ with $\lambda_{0}=\frac{\lambda c_{0}}{1-\lambda+\lambda c_{0}}$ and $c_{0}=(\frac{|{\rm supp}(G)|}{|{\rm supp}(F)|})^{1/n}$ such that up to null measure sets,
${\rm supp}(G)=c_{0}\,{\rm supp}(F)+x_{0},~~\text{and}~~{\rm supp}(H)=(1-\lambda+\lambda c_{0})\,{\rm supp}(F)+sx_{0},$
and for almost all $x\in{\rm supp}(F)$,
$\left\{\begin{aligned} &F(x)=\Phi(x)&&,\\ &G(c_{0}x+x_{0})=c_{0}^{\frac{1}{s}}\Phi(x)&&,\\ &H((1-\lambda+\lambda c_{0})x+\lambda x_{0})=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{0}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&.\end{aligned}\right$.
2.5. Fractional polar projection bodies
For $\alpha\in(-1,0)$ and measurable $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, the fractional polar projection body $\Pi^{*,-\alpha}f$ is defined by
$\rho_{\Pi^{*,-\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x)-f(x+ru)|dxdr.$
The affine fractional Sobolev inequalities (1.1) state that
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|\Pi^{*,-\alpha}f|^{\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy$
for $f\in W^{-\alpha,1}(\mathbb{R}^{n})$
Let $\mathcal{B}(\mathbb{R}^{n})$ denotes the class of Borel set in $\mathbb{R}^{n}$. Let $C_{c}^{\infty}(\mathbb{R}^{n};\mathbb{R}^{n})$ denote the set of smooth vector fields with compact support. Moreover, ${\rm div}\,\nu$ denotes the divergence of $\nu\in C_{c}^{\infty}(\mathbb{R}^{n};\mathbb{R}^{n})$. We say that $f\in L^{1}(\mathbb{R}^{n})$ is a function of bounded variation on $\mathbb{R}^{n}$ if there is a finite vector-valued Radon measure $\mathcal{D}f:\mathcal{B}(\mathbb{R}^{n})\rightarrow\mathbb{R}^{n}$ such that
$\int_{\mathbb{R}^{n}}\nu(x)\cdot n_{f}(x)d|\mathcal{D}f|(x)=-\int_{\mathbb{R}^{n}}f(x){\rm div}\,\nu(x)dx$
for every $\nu\in C_{c}^{\infty}(\mathbb{R}^{n};\mathbb{R}^{n})$, where $|\mathcal{D}f|:\mathcal{B}(\mathbb{R}^{n})\rightarrow[0,\infty)$ denotes the variation measure of $\mathcal{D}f$ and $n_{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ the Radon-Nikodym derivative of $\mathcal{D}f$ with respect to $|\mathcal{D}f|$.
For $f\in L^{1}(\mathbb{R}^{n})$ with bounded variation, the polar projection body $\Pi^{*}f$ is defined by
$\|u\|_{\Pi^{*}f}=\frac{1}{2}\int_{\mathbb{R}^{n}}|n_{f}(x)\cdot u|\,d|\mathcal{D}f|(x).$
For more information about $\Pi^{*}f$, we refer the reader to [44].
3. The star-shaped set ${\rm L}_{\alpha}f$
Building on the approach of Haddad and Ludwig [24, 25, 26], we introduce the following anisotropic integral. Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function. We consider
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy,$
where $K\subset\mathbb{R}^{n}$ is a star-shaped set with measurable radial function and $\alpha>0$.
Here, we first recall the definition of ${\rm L}_{\alpha}f$:
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr.$
Using Fubini’s theorem and polar coordinate, we obtain that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy$ $=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(y+z),f(y)\}}{\|z\|_{K}^{n-\alpha}}dzdy$
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\rho_{K}(ru)^{n-\alpha}r^{n-1}\min\{f(y),f(y+ru)\}drdudy$
$=\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(y),f(y+ru)\}dydrdu$
$=\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}du.$
That is,
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f).$ (3.1)
In particular,
$|{\rm L}_{n}f|$ $=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{{\rm L}_{n}f}(u)^{n}du$
$=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}r^{n-1}\min\{f(x),f(x+ru)\}dxdrdu$
$=\frac{1}{n}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy.$
Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function. Then the following hold:
For $\alpha\in(0,n)$, the dual mixed volume inequality (2.2) together with (3.1) implies
$\sup\Big\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\Big\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
For $\alpha>n$, the dual mixed volume inequality (2.3) together with (3.1) yields
$\inf\Big\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\Big\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
Moreover, in either regime a suitable dilate of ${\rm L}_{\alpha}f$ attains the supremum (respectively, infimum), provided $|{\rm L}_{\alpha}f|$ is finite.
We conclude this section with a result on the convexity of ${\rm L}_{\alpha}f$ when $f$ is log-concave.
Proposition 3.1.
Let $\alpha>0$. If $f:\mathbb{R}^{n}\to[0,\infty)$ is a log-concave, integrable function with compact support, then ${\rm L}_{\alpha}f$ is a convex body.
Proof.
For $u\in\mathbb{S}^{n-1}$, we define
$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx.$
Since $f$ is log-concave, Lemma 2.3 implies that $g$ is also log-concave.
Note that
$\rho_{{\rm L}_{\alpha}f}(u)=\Big(\int_{0}^{\infty}r^{\alpha-1}g(ru)dr\Big)^{1/\alpha}.$
It follows from [18, Corollary 4.2] that ${\rm L}_{\alpha}f$} is a convex body. ∎
4. Affine chord Sobolev inequalities
In this section, we establish the affine chord Sobolev inequalities stated in Theorem 1 and Theorem 2. To this end, we first prove a crucial lemma based on the Riesz rearrangement inequality.
Recall that a set $A\subset\mathbb{R}^{n}$ is homothetic to $B\subset\mathbb{R}^{n}$ if there exist $x\in\mathbb{R}^{n}$ and $a>0$ such that $A=x+aB$. Moreover, $A$ is said to be equivalent to $B$ if their symmetric difference $A\triangle B$ is a null set. The following lemma provides a general form of the Riesz rearrangement inequality; see also [10, 24, 25, 26].
Lemma 4.1.
Let $q>0$ and $K\subset\mathbb{R}^{n}$ be a star-shaped set with measurable radial function and $|K|>0$. For a measurable function $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ such that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy<\infty,$
there is
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{\|x-y\|_{K^{\star}}^{q}}dxdy.$
Equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost all $t>0$, the level set $\{f\geq t\}$ has measure zero or is homothetic to $D$ up to a set of measure zero.
Proof.
[…]
$\|z\|_{K}^{-q}=\int_{0}^{\infty}\operatorname{1}_{r^{-1/q}K}(z)dr.$
Using the Fubini’s theorem, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy$ $=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{\{f\geq t\}}(y)}{\|x-y\|_{K}^{q}}dxdydt$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{r^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq t\}}(y)dxdydrdt.$
By the Riesz rearrangement inequality, Theorem 2.2, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{r^{-1/q}K}$ $(x-y)\operatorname{1}_{\{f\geq t\}}(y)dxdy$ (4.1)
$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f^{\star}\geq t\}}(x)\operatorname{1}_{r^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f^{\star}\geq t\}}(y)dxdy,$
which implies that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{\|x-y\|_{K^{\star}}^{q}}dxdy.$
If equality holds, then (4.1) is an equality for almost all $(t,r)\in(0,\infty)^{2}$. For sufficiently large $r>0$, the assumptions of Theorem 2.2 are satisfied. Consequently, up to sets of measure zero, we have
$\{f\geq t\}=x+aD,\qquad r^{-1/q}K=bD,$
where $a,b>0$, $x\in\mathbb{R}^{n}$, and $D$ is a centered ellipsoid with the same volume as $K$. It follows that $K=\big(\frac{|K|}{|D|}\big)^{1/n}D$, showing that $D$ is independent of $t$ and $r$, which completes the proof. ∎
The next lemma shows that, for $\alpha\in(0,n)$, the volume of ${\rm L}_{\alpha}f$ does not increase under Schwarz symmetrals.
Lemma 4.2.
Let $\alpha\in(0,n)$. For non-negative $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$,
$|{\rm L}_{\alpha}f|\leq|{\rm L}_{\alpha}f^{\star}|.$
If $|{\rm L}_{\alpha}f^{\star}|<\infty$, there is equality if and only if $f$ is a translate of $f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
Proof.
We first assume that $|{\rm L}_{\alpha}f|<\infty$. By letting $q=n-\alpha$ in Lemma 4.1 and combining it with (3.1), we obtain
$\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f)\leq\tilde{V}_{\alpha}(K^{\star},{\rm L}_{\alpha}f^{\star}).$
With $K={\rm L}_{\alpha}f$ and the dual mixed volume inequality (2.2), we obtain
$|{\rm L}_{\alpha}f|$ $=\tilde{V}_{\alpha}({\rm L}_{\alpha}f,{\rm L}_{\alpha}f)$
$\leq\tilde{V}_{\alpha}(({\rm L}_{\alpha}f)^{\star},{\rm L}_{\alpha}f^{\star})$
$\leq|({\rm L}_{\alpha}f)^{\star}|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}$
$=|{\rm L}_{\alpha}f|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}},$
which implies that
$|{\rm L}_{\alpha}f|\leq|{\rm L}_{\alpha}f^{\star}|.$
The equality case follows from Lemma 4.1.
If $|{\rm L}_{\alpha}f|=\infty$, define
$f_{k}(x)=f(x)\operatorname{1}_{kB^{n}}(x),~~k\in\mathbb{N}.$
By the monotone convergence theorem,
$\lim_{k\rightarrow\infty}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f_{k}(x),f_{k}(x+ru)\}dxdr=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr,$
with the convergence being monotone increasing. Applying the monotone convergence theorem once more then gives
$\lim_{k\rightarrow\infty}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{$ $f_{k}(x),f_{k}(x+ru)\}dxdr\Big)^{\frac{n}{\alpha}}du$
$=\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr\Big)^{\frac{n}{\alpha}}du.$
That is,
$\lim_{k\rightarrow\infty}|{\rm L}_{\alpha}f_{k}|=|{\rm L}_{\alpha}f|=\infty.$ (4.2)
Since $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$ and $f_{k}$ has compact support, $|{\rm L}_{\alpha}f_{k}|<\infty$. By the definition of the Schwarz symmetral, we have $(f_{k})^{\star}\leq f^{\star}$. Then the first part of the proof implies that
$|{\rm L}_{\alpha}f_{k}|\leq|{\rm L}_{\alpha}(f_{k})^{\star}|\leq|{\rm L}_{\alpha}f^{\star}|.$
Taking the limit as $k\rightarrow\infty$ in the previous inequality, we obtain from (4.2) that
$|{\rm L}_{\alpha}f^{\star}|=\infty,$
which gives the desired result. ∎
Proof of Theorem 1.
By (3.1) and the dual mixed volume inequality (2.2), we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f)\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
There is equality precisely when ${\rm L}_{\alpha}f$ is a ball, which is the case for radially symmetric functions.
For the second inequality, noting that $\|f\|_{\frac{n}{n+\alpha}}=\|f^{\star}\|_{\frac{n}{n+\alpha}}$ and that ${\rm L}_{\alpha}f^{\star}$ is a ball, we obtain from the chord Sobolev inequality (1.3), (3.1), and Lemma 4.2 that
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}$ $\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy$
$=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f^{\star})$
$=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}$
$\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
If equality holds throughout, it follows from the chord Sobolev inequality that $f^{\star}=c\operatorname{1}_{B}$, where $B$ is a ball. By Lemma 4.2, we have $f=f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. This characterizes the equality cases. ∎
For $\alpha>n$, we state the following analogue of Lemma 4.1, which also generalizes [4, Lemma 3.2]. The proof follows a similar approach and is included for completeness.
Lemma 4.3.
Let $q>0$ and $K\subset\mathbb{R}^{n}$ be a star-shaped set with measurable radial function and $|K|>0$. If $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is an integrable function such that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}\|x-y\|_{K}^{q}dxdy<\infty,$
then
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}\|x-y\|_{K}^{q}dxdy\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f^{\star}(x),f^{\star}(y)\}\|x-y\|_{K^{\star}}^{q}dxdy.$
Equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost $t>0$, the level set $\{f\geq t\}$ has measure zero or is homothetic to $D$ up to sets of measure zero.
Proof.
[…]
$\|z\|_{K}^{q}=\int_{0}^{\infty}\operatorname{1}_{\mathbb{R}^{n}\setminus r^{1/q}K}(z)dr.$
Using Fubini’s theorem, we obtain
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),$ $f(y)\}\|x-y\|_{K}^{q}dxdy$ (4.3)
$=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\|x-y\|_{K}^{q}\operatorname{1}_{\{f\geq t\}}(y)dxdydt$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{\mathbb{R}^{n}\setminus r^{1/q}K}(x-y)\operatorname{1}_{\{f\geq t\}}(y)dxdydtdr$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)(1-\operatorname{1}_{r^{1/q}K}(x-y))\operatorname{1}_{\{f\geq t\}}(y)dxdydtdr.$
The Riesz rearrangement inequality, Theorem 2.2, implies that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{r^{1/q}K}(x-y)$ $\operatorname{1}_{\{f\geq t\}}(y)dxdy$ (4.4)
$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f^{\star}\geq t\}}(x)\operatorname{1}_{r^{1/q}K^{\star}}(x-y)\operatorname{1}_{\{f^{\star}\geq t\}}(y)dxdy.$
Note that $\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)dx<\infty$ for almost all $t>0$, as $f\in L^{1}(\mathbb{R}^{n})$. Combing (4.4) with the calculation in (4.3), we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}\|x-y\|_{K}^{q}dxdy\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f^{\star}(x),f^{\star}(y)\}\|x-y\|_{K^{\star}}^{q}dxdy.$ (4.5)
Moreover, if equality holds in (4.5), then there exists a null set $N\subset(0,\infty)^{2}$ such that equality holds in (4.4) for all $(t,r)\in(0,\infty)^{2}\setminus N$. For almost all $r\in(0,\infty)$, we have $(t,r)\in(0,\infty)^{2}\setminus N$ for almost all $t>0$. For $r$ sufficiently small, Theorem 2.2 guarantees the existence of a centered ellipsoid $D$ and $x\in\mathbb{R}^{n}$ such that
$\{f\geq t\}=x+aD\quad\text{and }\quad r^{1/q}K=bD.$
Since $K=(\frac{|K|}{|D|})^{1/n}D$, the ellipsoid $D$ is independent of $r$ and $t$, which concludes the proof. ∎
Similarly, for $\alpha>n$, we show that $|{\rm L}_{\alpha}f|$ does not decrease under Schwarz symmetrals.
Lemma 4.4.
Let $\alpha>n$ and $f:\mathbb{R}^{n}\to[0,\infty)$ be integrable. If $|{\rm L}_{\alpha}f|<\infty$, then
$|{\rm L}_{\alpha}f|\geq|{\rm L}_{\alpha}f^{\star}|.$
Equality holds if and only if $f$ is a translate of $f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
Proof.
By letting $q=\alpha-n$ in Lemma 4.3 and combining it with (3.1), we obtain
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f)\geq n\tilde{V}_{\alpha}(K^{\star},{\rm L}_{\alpha}f^{\star}).$
With $K={\rm L}_{\alpha}f$ and the dual mixed volume inequality (2.3), we have
$|{\rm L}_{\alpha}f|$ $=\tilde{V}_{\alpha}({\rm L}_{\alpha}f,{\rm L}_{\alpha}f)$
$\geq\tilde{V}_{\alpha}(({\rm L}_{\alpha}f)^{\star},{\rm L}_{\alpha}f^{\star})$
$\geq|({\rm L}_{\alpha}f)^{\star}|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}$
$=|{\rm L}_{\alpha}f|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}},$
which implies that
$|{\rm L}_{\alpha}f|\geq|{\rm L}_{\alpha}f^{\star}|.$
The equality cases follows from Lemma 4.3. ∎
Building on the preceding lemmas, we now establish the affine chord Sobolev inequalities for $\alpha>n$.
Proof of Theorem 2.
If the right-most side is $\infty$, the inequality holds trivially. We then assume that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy<\infty.$
It follows from (3.1) and the dual mixed volume inequality (2.3) that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f)\geq n\omega^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
By the equality case in the dual mixed volume inequality, ${\rm L}_{\alpha}f$ must be a ball, which occurs when $f$ is radially symmetric.
For the first inequality, we may assume that $|{\rm L}_{\alpha}f|$ is finite. Note that ${\rm L}_{\alpha}f^{\star}$ is a ball. Then Lemma 4.4 and (3.1) imply that
$n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f^{\star})=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy.$
Combining the chord Sobolev inequality (1.4) with $\|f\|_{1}=\|f^{\star}\|_{1}$ and $\|f\|_{\infty}=\|f^{\star}\|_{\infty}$, we deduce
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
The equality case follows from the optimizers of (1.4) together with Lemma 4.4. ∎
5. Radial mean bodies for log-concave functions
In this section, we begin by recalling some basic notation and properties of the classical radial mean bodies. Let $K\subset\mathbb{R}^{n}$ be a convex body. For $\alpha>-1$ and $u\in\mathbb{S}^{n-1}$, Gardner and Zhang [18] defined the radial $\alpha$-th mean body of $K$ by its radial function:
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{1}{|K|}\int_{K}\rho_{K-x}(u)^{\alpha}dx$
for $\alpha\neq 0$ and by
$\log\rho_{{\rm R}_{0}K}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx.$
They further defined ${\rm R}_{\infty}K$ consistently by
$\rho_{{\rm R}_{\infty}K}(u)=\lim_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}K}(u)=\max_{x\in K}\rho_{K-x}(u),$
so that ${\rm R}_{\infty}K$ coincides with the difference body of $K$, namely,
${\rm R}_{\infty}K={\rm D}K=K-K.$
5.1. Affine isoperimetric inequalities for $\boldsymbol{{\rm R}_{\alpha}f}$
Haddad and Ludwig [24] proved the following affine isoperimetric inequality for ${\rm R}_{\alpha}K$, with full equality characterization.
Theorem 5.1.
[24, Theorem 20] Let $K\subset\mathbb{R}^{n}$ be a convex body. Then
$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha\in(-1,n),$
$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha>n,$
with equality if and only if $K$ is an ellipsoid.
For $\alpha>-1$, the authors [4] defined the radial mean body ${\rm R}_{\alpha}f$ for a non-zero, log-concave function $f\in L^{1}(\mathbb{R}^{n})$ by
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t),~~\alpha\neq 0,$ (5.1)
and
$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq t\}}(u)d\mu_{f}(t),$ (5.2)
where $d\mu_{f}(t)=\frac{|\{f\geq t\}|}{\|f\|_{1}}dt$ is a probability measure. The following result in [4] shows that ${\rm R}_{\alpha}f$ is proportional to the fractional polar projection body $\Pi^{*,-\alpha}f$ and the star-shaped set ${\rm L}_{\alpha}f$. Moreover, by the continuity of $\rho_{{\rm R}_{\alpha}f}(u)$ with respect to $\alpha$, the authors also gave an explicit formula for $\rho_{{\rm R}_{0}f}$.
Lemma 5.2.
[4, Proposition 6.2, Lemma 6.4] For non-zero, log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x)-f(x+ru)|dxdr$ (5.3)
for $\alpha\in(-1,0)$ and
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{\alpha}{\|f\|_{1}}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr$ (5.4)
for $\alpha>0$.
For $\alpha=0$,
$\log\rho_{{\rm R}_{0}f}(u)=-\gamma+\int_{0}^{\infty}\frac{1}{r}\Big(\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx-e^{-r}\Big)dr,$
where $\gamma=\Gamma^{\prime}(1)$ is the Euler constant.
The following lemma describes the affine covariance of ${\rm R}_{\alpha}f$, where ${\rm GL}(n)$ denotes the group of general linear transformations on $\mathbb{R}^{n}$.
Lemma 5.3.
Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
${\rm R}_{\alpha}(cf)={\rm R}_{\alpha}f~~\text{and}~~{\rm R}_{\alpha}(f\circ\tau_{y})={\rm R}_{\alpha}f,$
where $c>0$ is a constant, $y\in\mathbb{R}^{n}$ and $f\circ\tau_{y}(x)=f(x-y)$. Moreover, for $\phi\in{\rm GL}(n)$,
${\rm R}_{\alpha}(f\circ\phi^{-1})=\phi{\rm R}_{\alpha}f.$
Proof.
By (5.3) and (5.4), the radial mean body ${\rm R}_{\alpha}f$ is invariant under scaling and translations. It suffices to show ${\rm R}_{\alpha}(f\circ\phi^{-1})=\phi{\rm R}_{\alpha}f$.
We treat the case $\alpha>n$. The case $\alpha\in(-1,0)$ can be handled similarly. For $\phi\in{\rm GL}(n)$, note that $\|f\circ\phi^{-1}\|_{1}=|\det\phi|\|f\|_{1}$. Since
$\int_{\mathbb{R}^{n}}\min\{f(\phi^{-1}x),f(\phi^{-1}(x+ru))\}dx=|\det\phi|\int_{\mathbb{R}^{n}}\min\{f(x),f(x+r\phi^{-1}u)\}dx,$
by (5.4), we have
$\rho_{{\rm R}_{\alpha}(f\circ\phi^{-1})}(u)^{\alpha}=\rho_{{\rm R}_{\alpha}f}(\phi^{-1}u)^{\alpha}=\rho_{\phi{\rm R}_{\alpha}f}(u)^{\alpha},$
which concludes the proof. ∎
Recall that the affine chord Sobolev inequalities can be written as
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}$ $\leq n\omega_{n}^{\frac{n-\alpha}{n}}|\Pi^{*,-\alpha}f|^{\frac{\alpha}{n}},$ $\alpha\in(-1,0),$ (5.5)
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}$ $\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}},$ $\alpha\in(0,n)$
and
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}},~~\alpha>n.$ (5.6)
By (5.4) and the affine chord Sobolev inequalities, we establish the following affine isoperimetric inequalities for ${\rm R}_{\alpha}f$. We first focus on non-negative $\alpha$.
Note that, when $\alpha=n$ and $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ is log-concave, the chord Sobolev inequalities become the following two trivial inequalities:
$\frac{\|f\|_{1}^{2}}{\|f\|_{\infty}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy\leq\|f\|_{\frac{1}{2}},$ (5.7)
where equality holds if and only if $f=c\operatorname{1}_{K}$ for some convex body $K\subset\mathbb{R}^{n}$ and constant $c>0$. Since
$|{\rm R}_{n}f|=\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy,$
and $|{\rm R}_{n}K|=|K|$ for any convex body $K\subset\mathbb{R}^{n}$, both (5.8) and (5.9) hold for $\alpha=n$ and equality holds if and only if $f=c\operatorname{1}_{K}$.
Theorem 5.4.
Let $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ be log-concave and non-zero. Then
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},\quad\alpha\in(-1,0)\cup(0,n),$ (5.8)
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{1}\|f\|_{\infty}^{-1}}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},\quad\alpha>n.$ (5.9)
In both cases, equality holds if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid.
Proof.
Inequality (5.9) and the equality characterization follow from (5.6) directly. We turn to $\alpha\in(-1,0)\cup(0,n)$. By the affine chord Sobolev inequality (5.5), we obtain
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{n}{n+\alpha}}^{n/\alpha}\|f\|_{1}^{-n/\alpha}}\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|}.$
For brevity, we denote by $\bar{f}(x)=f(x)/\|f\|_{1}$. Then $\bar{f}(x)dx$ is a probability measure on $\mathbb{R}^{n}$. Combining this with the monotonicity of $L^{p}$ norms, for $\alpha\in(-1,n)\setminus\{0\}$, we have
$\|f\|_{\frac{n}{n+\alpha}}^{n/\alpha}\|f\|_{1}^{-n/\alpha}$ $=\Big(\int_{\mathbb{R}^{n}}\bar{f}(x)^{\frac{n}{n+\alpha}}dx\Big)^{\frac{n+\alpha}{\alpha}}$ (5.10)
$=\Big(\int_{\mathbb{R}^{n}}\big(\bar{f}(x)^{-1}\big)^{\frac{\alpha}{n+\alpha}}\bar{f}(x)dx\Big)^{\frac{n+\alpha}{\alpha}}\leq\Big(\int_{\mathbb{R}^{n}}\bar{f}(x)^{\frac{1}{2}}dx\Big)^{2}=\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1},$
where equality holds if and only if $f=c\operatorname{1}_{K}$ for some $c>0$ and convex body $K\subset\mathbb{R}^{n}$. Therefore, we obtain
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|}.$ (5.11)
If equality holds, then (5.10) implies that $f=c\operatorname{1}_{K}$ where $K\subset\mathbb{R}^{n}$ is a convex body and $c>0$. Hence (5.11) reduces to the geometric inequality (5.1), whose equality cases imply that $K$ is an ellipsoid. ∎
Next we consider the case $\alpha=0$. For non-zero, log-concave $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ and fixed $u\in\mathbb{S}^{n-1}$, the radial function $\rho_{{\rm R}_{\alpha}f}(u)$ is continuous in $\alpha>-1$, as shown in [4]. Therefore, taking limits on both sides of (5.11) yields the affine isoperimetric inequality for ${\rm R}_{0}f$, but the equality characterization does not follow from this argument. In the following theorem, we use (5.1) together with the definition of ${\rm R}_{0}f$ to establish the inequality with its equality characterization.
Theorem 5.5.
Let $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
$\frac{|{\rm R}_{0}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}\leq\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|},$
where equality holds if and only if $f$ is the characteristic function of an ellipsoid.
Proof.
By (5.2),
$\rho_{{\rm R}_{0}f}(u)^{n}=\exp\Big(\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq t\}}(u)^{n}d\mu_{f}(t)\Big).$
Since $e^{t}$ is a convex function on $\mathbb{R}$ and $\mu_{f}$ is a probability on $\mathbb{R}$, it follows from Jensen’s inequality that
$\rho_{{\rm R}_{0}f}(u)^{n}\leq\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{0}\{f\geq t\}}(u)^{n}d\mu_{f}(t).$
This, together with (5.1), implies that
$|{\rm R}_{0}f|=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{{\rm R}_{0}f}(u)^{n}du$ $\leq\frac{1}{n}\int_{\mathbb{S}^{n-1}}\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{0}\{f\geq t\}}(u)^{n}d\mu_{f}(t)du$
$=\int_{0}^{\|f\|_{\infty}}|{\rm R}_{0}\{f\geq t\}|d\mu_{f}(t)$
$\leq\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|}\int_{0}^{\|f\|_{\infty}}\frac{|\{f\geq t\}|^{2}}{\|f\|_{1}}dt$
$=\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|f\|_{1}}dxdy.$
By (5.7), we obtain
$\frac{|{\rm R}_{0}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}\leq\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|}.$ (5.12)
If equality holds, then
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy=\|f\|_{\frac{1}{2}},$
which implies that $f=c\operatorname{1}_{K}$ for some convex body $K\subset\mathbb{R}^{n}$ and $c>0$. Hence (5.12) reduces to the geometric inequality in (5.1). By the equality characterization of (5.1), $K$ is an ellipsoid. ∎
5.2. Monotonicity of $\boldsymbol{{\rm R}_{\alpha}f}$
To consider the monotonicity of radial mean bodies, we recall the following result by Haddad and Ludwig [24].
Lemma 5.6.
[24, Theorem 3] For $\alpha\in(-1,0)$ and $f\in L^{1}(\mathbb{R}^{n})$ with bounded variation,
$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|f\|_{1}\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\rho_{\Pi^{*}f}(u)^{-1}$
for $u\in\mathbb{S}^{n-1}$.
We mainly use Lemma 5.6 for log-concave functions $f\in L^{1}(\mathbb{R}^{n})$. By [4, Lemma 5.4], $\mathcal{H}^{n-1}(\partial\{f\geq t\})$ is a integrable function on $\mathbb{R}$. Hence the coarea formula [14, Theorem 5.9], implies that $f$ has bounded variation. The opposite endpoint $\alpha=\infty$ will be considered below.
Lemma 5.7.
Let $f\in L^{1}(\mathbb{R}^{n})$ be log-concave and non-zero. If $f$ has compact support, then
$\lim_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)=\sup_{t>0}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}\,{\rm supp}(f)}(u),$
for $u\in\mathbb{S}^{n-1}$.
Proof.
Since Jensen’s inequality implies that ${\rm R}_{\alpha}K\subset{\rm D}K$ for $\alpha>-1$, we have
$\rho_{{\rm R}_{\alpha}f}(u)\leq\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t)\Big)^{1/\alpha}.$
Letting $\alpha\rightarrow\infty$, we obtain
$\limsup_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\leq\sup_{t>0}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}\,{\rm supp}(f)}(u).$ (5.13)
On the other hand, by the reversed inclusion ${\rm D}K\subset c_{n,\alpha}{\rm R}_{\alpha}K$ in (1.12),
$\rho_{{\rm R}_{\alpha}f}(u)\geq c_{n,\alpha}^{-1}\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t)\Big)^{1/\alpha},$
where $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$. Note that $\lim_{\alpha\rightarrow\infty}c_{n,\alpha}=1$. Therefore we have
$\liminf_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\geq\sup_{t>0}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}\,{\rm supp}(f)}(u).$ (5.14)
Combining (5.13) with (5.14), we then conclude the proof. ∎
This yields the natural definition of ${\rm R}_{\infty}f$ given below.
Definition 5.8.
Let $f\in L^{1}(\mathbb{R}^{n})$ be log-concave with compact support. The body ${\rm R}_{\infty}f$ is given by
${\rm R}_{\infty}f={\rm D}\,{\rm supp}(f).$
If $f$ is not compactly supported, ${\rm R}_{\infty}f=\mathbb{R}^{n}$.
Recall that in the geometric case, Jensen’s inequality implies that
${\rm R}_{\alpha}K\subset{\rm R}_{\beta}K\subset{\rm D}K,$ (5.15)
where Gardner and Zhang [18] further showed that
$\Pi^{*}K\subset((1+\alpha)|K|)^{1/\alpha}{\rm R}_{\alpha}K\subset((1+\beta)|K|)^{1/\beta}{\rm R}_{\beta}K\subset{\rm D}K$ (5.16)
for $1<\alpha<\beta$.
Applying Jensen’s inequality, together with (5.15), we immediately obtain
${\rm R}_{\alpha}f\subset{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$
for $1<\alpha<\beta<\infty$. The following theorem provides a functional analogue of (5.16).
Theorem 5.9.
For non-zero, log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\|f\|_{\infty}\Pi^{*}f\subset\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$
for $1<\alpha<\beta<\infty$.
Proof.
If $f$ is not compactly supported, the right-most inclusion is trivial since ${\rm R}_{\infty}f=\mathbb{R}^{n}$. We now assume that $f$ has compact support.
By the definition of ${\rm R}_{\alpha}f$ in (5.1),
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}(1+\alpha)|\{f\geq t\}|\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}dt\Big)^{1/\alpha}$
$=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\alpha,t}{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}dt\Big)^{1/\alpha},$
where $l_{\alpha,t}=((1+\alpha)|\{f\geq t\}|)^{1/\alpha}$. It follows from (5.16) that
$l_{\alpha,t}\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)\leq l_{\beta,t}\rho_{{\rm R}_{\beta}\{f\geq t\}}(u),$
for $1<\alpha<\beta$.
Jensen’s inequality then implies that
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,t}{\rm R}_{\beta}\{f\geq t\}}(u)^{\alpha}dt\Big)^{1/\alpha}$
$\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,t}{\rm R}_{\beta}\{f\geq t\}}(u)^{\beta}dt\Big)^{1/\beta}$
$=\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}\rho_{{\rm R}_{\beta}f}(u),$
which gives
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f.$
Letting $\beta\rightarrow\infty$, together with $\lim_{\beta\rightarrow\infty}(1+\beta)^{1/\beta}=1$, we obtain the right-most inclusion. Letting $\alpha\rightarrow-1^{+}$, the left-most inclusion follows from Lemma 5.6. ∎
6. Reverse inclusion for radial mean bodies
In the previous section, we recalled the monotonicity relations (5.15) and (5.16) for radial mean bodies and extended them to their functional analogues. In this section, we turn to the reverse inclusion, which is considerably more delicate. We begin by stating the following theorem established by Gardner and Zhang [18].
Theorem 6.1.
[18, Theorem 5.5] Let $K$ be a convex body in $\mathbb{R}^{n}$. If $1<\alpha<\beta$, then
${\rm D}K\subset c_{n,\beta}{\rm R}_{\beta}K\subset c_{n,\alpha}{\rm R}_{\alpha}K\subset n|K|\Pi^{*}K.$
In each inclusion equality holds if and only if $K$ is a simplex.
When $\beta=n$, the left-most inclusion implies the Rogers-Shephard inequality:
$|{\rm D}K|\leq\binom{2n}{n}|K|.$
When $\alpha=n$, the right-most inclusion yields Zhang’s inequality:
$n^{-n}\binom{2n}{n}=|\triangle|^{n-1}|\Pi^{*}\triangle|\leq|K|^{n-1}|\Pi^{*}K|,$
where $\triangle$ denotes a simplex in $\mathbb{R}^{n}$.
To extend Theorem 6.1 to the functional setting, we rely on the following result from [26]. It generalizes the classical result of Marshall, Olkin and Proschan [38], whose proof was later simplified by Milman and Pajor [40].
Lemma 6.2.
[26, Lemma 17] Let $\omega:[0,\infty)\to[0,\infty)$ be decreasing with
$0<\int_{0}^{\infty}r^{\alpha-1}\omega(r)dr<\infty$
for $\alpha>0$ and
$0<\int_{0}^{\infty}r^{\alpha-1}(\omega(0)-\omega(r))dr<\infty$
for $\alpha\in(-1,0)$. If $\varphi:[0,\infty)\to[0,\infty)$ is non-zero, with $\varphi(0)=0$, and such that $r\mapsto\varphi(r)$ and $r\mapsto{\varphi(r)}/{r}$ are increasing on $(0,\infty)$, then
$\zeta(\alpha)=\left\{\begin{aligned} &\left(\frac{\int_{0}^{\infty}r^{\alpha-1}\omega(\varphi(r))dr}{\int_{0}^{\infty}r^{\alpha-1}\omega(r)dr}\right)^{1/\alpha}&&\text{for}~~\alpha>0,\\ &\exp\left(\int_{0}^{\infty}\frac{\omega(\varphi(r))-\omega(r)}{r\omega(0)}dr\right)&&\text{for}~~\alpha=0,\\ &\left(\frac{\int_{0}^{\infty}r^{\alpha-1}(\omega(\varphi(r))-\omega(0))dr}{\int_{0}^{\alpha-1}(\omega(r)-\omega(0))dr}\right)^{1/\alpha}&&\text{for}~~-1<\alpha<0,\end{aligned}\right$.
is a continuous, decreasing function of $\alpha$ on $(-1,\infty)$. Moreover, $\zeta$ is constant on $(-1,\infty)$ if $\varphi(r)=\lambda r$ on $0,\infty)$ for some $\lambda>0$.
In particular, taking $\omega(r)=e^{-r}$ and $\omega(r)=(1-sr)_{+}^{1/s}$ with $s>0$ yields the classical analytic continuation formulas for the Gamma and Beta functions,
[…]
and
[…]
Under the assumptions of Lemma 6.2 with strictly decreasing $\omega(r)$, if $\zeta(\alpha)$ is constant on $(-1,\infty)$, say $\zeta(\alpha)=\lambda$, then $\varphi(r)=\lambda r$ for all $r\geq 0$. The result follows directly from Milman and Pajor [40], but we provide a proof for completeness.
Lemma 6.3.
Let $\omega$, $\varphi$, and $\zeta$ be as in Lemma 6.2. Suppose $\omega$ is strictly decreasing and $1<\alpha<\beta<\infty$. Then $\zeta(\alpha)=\zeta(\beta)$ implies $\varphi(r)=\lambda r$ with $\lambda=\zeta(\alpha)^{-1}$.
Proof.
Let $\lambda=\zeta(\alpha)^{-1}$. We first consider $0<\alpha<\beta<\infty$. For $\alpha>0$, we then have
$\int_{0}^{\infty}\omega(\lambda r)r^{\alpha-1}dr=\int_{0}^{\infty}\omega(\varphi(r))r^{\alpha-1}dr.$
We set $\Psi(t)=\int_{t}^{\infty}(\omega(\lambda r)-\omega(\varphi(r)))dr$, where $\Psi(0)=\Psi(\infty)=0$. Since $\varphi(r)/r$ is increasing while $\omega$ is decreasing, we have
$\omega(\varphi(r))>\omega(\lambda r),~~\text{if}~~{\varphi(r)}/{r}<\lambda,$
and
$\omega(\varphi(r))\leq\omega(\lambda r),~~\text{if}~~{\varphi(r)}/{r}\geq\lambda.$
This implies the sign of $\omega(\lambda t)-\omega(\varphi(t))=-\Psi^{\prime}(t)$. Therefore, $\Psi$ is first increasing and then decreasing, which yields that $\Psi(t)\geq 0$ for all $t\geq 0$. That is
$\int_{t}^{\infty}\omega(\lambda r)r^{\alpha-1}dr\geq\int_{t}^{\infty}\omega(\varphi(r))r^{\alpha-1}dr.$
For $\beta>\alpha>0$, we have
$\int_{0}^{\infty}\omega(\varphi(r))r^{\beta-1}dr$ $=(\beta-\alpha)\int_{0}^{\infty}t^{\beta-\alpha-1}\Big(\int_{t}^{\infty}\omega(\varphi(r))r^{\alpha-1}dr\Big)dt$
$\leq(\beta-\alpha)\int_{0}^{\infty}t^{\beta-\alpha-1}\Big(\int_{t}^{\infty}\omega(\lambda r)r^{\alpha-1}dr\Big)dt$
$=\int_{0}^{\infty}\omega(\lambda r)r^{\beta-1}dr=\frac{1}{\lambda^{\beta}}\int_{0}^{\infty}\omega(r)r^{\beta-1}dr,$
which implies $\zeta(\alpha)\geq\zeta(\beta)$. If $\zeta(\alpha)=\zeta(\beta)=\lambda^{-1}$, equality holds in the inequality above, which yields that $\Psi(t)=0$ for all $t\geq 0$. Then we have
$\Psi^{\prime}(t)=\omega(\varphi(t))-\omega(\lambda t)=0.$
Since $\omega$ is strictly decreasing, we obtain $\varphi(t)=\lambda t$.
For $1<\alpha<\beta<0$, the proof is the same. Then we consider $1<\alpha<0<\beta<\infty$. Since $\zeta$ is decreasing, we have
$\zeta(\beta)\leq\zeta(0)\leq\zeta(\alpha).$
As assumed $\zeta(\alpha)=\zeta(\beta)=\lambda^{-1}$, we obtain that $\zeta$ is a constant on $(0,\beta)$, by the first case, we have $\varphi(t)=\lambda t$. ∎
6.1. Results for log-concave functions
In this subsection, for the sake of completeness, we provide a proof of (1.13), which extends (1.12) to the class of log-concave functions and was recently established in [28, Theorem 1.6]. The following result from [3] serves as a main tool in the equality characterization of (1.13).
Lemma 6.4.
[3, Lemma 4.2] Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave such that $f(o)=\|f\|_{\infty}$. Let
$g(x)=\int_{\mathbb{R}^{n}}\min\{f(y),f(y+x)\}dy.$
Then for every $x\in\mathbb{R}^{n}$ and $\lambda\in(0,1)$, $g(\lambda x)=g(o)^{1-\lambda}g(x)^{\lambda}$ if and only if $f(x)=\|f\|_{\infty}e^{-\|x\|_{\triangle}}$ with $\triangle$ a simplex containing the origin.
For the reader’s convenience, we restate (1.13) below and give a proof.
Theorem 6.5.
Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f\subset 2\|f\|_{1}\Pi^{*}f$ (6.2)
for $1<\alpha<\beta<\infty$. In each inclusion, there is equality if and only if $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ where $a=\|f\|_{\infty}$ and $\triangle$ is a simplex containing $x_{0}$.
Proof.
By Lemma 2.3,
$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$
is an even, log-concave function and attains maximum at $r=0$.
For fixed $u\in\mathbb{S}^{n-1}$, we apply Lemma 6.2 to $\omega(r)=g(o)e^{-r}$ and $\varphi(r)=-\log(g(ru)/g(o))$. We observe that
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}g(ru)dr}{\int_{0}^{\infty}r^{\alpha-1}g(o)e^{-r}dr}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
for $\alpha>0$.
For $\alpha\in(-1,0)$, note that
$g(o)-g(ru)$ $=\|f\|_{1}-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$ (6.3)
$=\frac{1}{2}\int_{\mathbb{R}^{n}}f(x)dx+\frac{1}{2}\int_{\mathbb{R}^{n}}f(x+ru)dx-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$
$=\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x)-f(x+ru))_{+}dx+\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x+ru)-f(x))_{+}dx$
$=\frac{1}{2}\int_{\mathbb{R}^{n}}|f(x)-f(x+ru)|dx,$
where $h(x)_{+}$ denotes the positive part of the function $h(x)$. Then we have
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}(g(o)-g(ru))dr}{\int_{0}^{\infty}r^{\alpha-1}g(o)(1-e^{-r})dr}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
for $\alpha\in(-1,0)$.
Therefore Lemma 6.2 implies that
$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f.$ (6.4)
Letting $\alpha\rightarrow-1$, Lemma 5.6 gives the right inclusion. Since $\omega(r)=e^{-r}$ is strictly decreasing, by Lemma 6.3, equality holds in (6.4) if and only if $g(ru)=g(o)e^{-\lambda r}$ for some $\lambda>0$. Then Lemma 6.4 shows that $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for $a=\|f\|_{\infty}$ and $x_{0}\in\triangle$, where $\triangle$ is a simplex containing the origin. ∎
If we set $f(x)=e^{-\|x\|_{K}}$ for a convex body $K\subset\mathbb{R}^{n}$, then inclusion (6.2) in Theorem 6.5 reduces to the geometric case stated in Theorem 6.1. However, in this functional extension the endpoint ${\rm R}_{\infty}f$ is $\mathbb{R}^{n}$, so the inclusion no longer extends to that limit.
This observation motivates the consideration of the embedding $K\mapsto\operatorname{1}_{K}$ in studying functional extensions of Theorem 6.1. The following counterexample shows, however, that this extension fails for general log-concave functions, which naturally leads to the study of $s$-concave functions in the next subsection.
Example 6.6.
Let $n=1$ and $f(x)=2x\operatorname{1}_{[0,1]}(x)$. It is clear that $\log f(x)=\log x+\log 2$ for $x\in[0,1]$, which is concave on ${\rm supp}(f)$. Then we have
$g(r)=(1-|r|)_{+}^{2}.$
In this case, we consider $\beta=1$ and $\alpha=\frac{1}{2}$ and obtain that
$c_{1,1}{\rm R}_{1}f=\Big[-\frac{2}{3},\frac{2}{3}\Big],\quad c_{1,\frac{1}{2}}{\rm R}_{\frac{1}{2}}f=\Big[-\frac{16}{25},\frac{16}{{25}}\Big],$
which is not consistent with the desired result $c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f$.
6.2. Results for $s$-concave functions
In this subsection, we focus on the case of $s$-concave functions with $s>0$. We continue to use the notation
$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx.$
The following lemma is a direct corollary from the Borell-Brascamp-Lieb inequality, which is also mentioned in [17, Theorem 11.2].
Lemma 6.7.
Let $s>0$ and $u\in\mathbb{S}^{n-1}$ be fixed. For an $s$-concave, integrable function $f$, the function $r\mapsto g(ru)$ is $s/(ns+1)$-concave on $\mathbb{R}$.
Proof.
We denote by
$N_{f}(x,r)=\min\{f(x),f(x+ru)\}.$
First, we prove that $N_{f}(x,r)$ is $s$-concave on $\mathbb{R}^{n}\times\mathbb{R}$. For $x_{0},x_{1}\in\mathbb{R}^{n}$, $r_{0},r_{1}\in\mathbb{R}$ and $\lambda\in[0,1]$, we denote $x_{\lambda}=(1-\lambda)x_{0}+\lambda x_{1}$ and $r_{\lambda}=(1-\lambda)r_{0}+\lambda r_{1}$. Note that
$N_{f}(x_{\lambda},r_{\lambda})^{s}=\min\{f(x_{\lambda})^{s},f(x_{\lambda}+r_{\lambda}u)^{s}\}.$
Since $f$ is $s$-concave and $\min\{\cdot,\cdot\}$ is a concave function on $\mathbb{R}\times\mathbb{R}$, we have
$N_{f}(x_{\lambda},r_{\lambda})^{s}$ $\geq\min\{(1-\lambda)f(x_{0})^{s}+\lambda f(x_{1})^{s},(1-\lambda)f(x_{0}+r_{0}u)^{s}+\lambda f(x_{1}+r_{1}u)^{s}\}$
$\geq(1-\lambda)\min\{f(x_{0})^{s},f(x_{0}+r_{0}u)^{s}\}+\lambda\min\{f(x_{1})^{s},f(x_{1}+r_{1}u)^{s}\}$
$=(1-\lambda)N_{f}(x_{0},r_{0})^{s}+\lambda N_{f}(x_{1},r_{1})^{s},$
which implies the $s$-concavity of $N_{f}$.
For $\lambda\in[0,1]$, and fixed $r_{0},r_{1}\in\mathbb{R}$, we set
$F(x)=N_{f}(x,r_{0}),~~G(x)=N_{f}(x,r_{1}),~~H(x)=N_{f}(x,r_{\lambda}),$
where $r_{\lambda}=(1-\lambda)r_{0}+\lambda r_{1}$. By the $s$-concavity of $N_{f}$, we have
$H((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(F(x),G(y)),$
where $\mathcal{M}_{\lambda}^{s}$ denotes the $s$-mean (see Section 2.4). The Borell-Brascamp-Lieb inequality (2.4) then implies that
$\int_{\mathbb{R}^{n}}H\geq\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}F,\int_{\mathbb{R}^{n}}G\Big),$ (6.5)
that is, the $\frac{s}{ns+1}$-concavity of $g(ru)$ in $r$. ∎
We begin the proof of Theorem 3 by stating the inclusion result as the following theorem, the equality case will be addressed separately.
Theorem 6.8.
Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero $s$-concave function with compact support. If $1<\alpha<\beta$, then
${\rm R}_{\infty}f\subset c_{n,\beta}(s){\rm R}_{\beta}f\subset c_{n,\alpha}(s){\rm R}_{\alpha}f\subset(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$ (6.6)
where $c_{n,\alpha}(s)=((n+\frac{1}{s})B(\alpha+1,n+\frac{1}{s}))^{-1/\alpha}$.
Proof.
By Lemma 6.7, the function $g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$ is $s^{\prime}$-concave with $s^{\prime}=s/(ns+1)$. We apply Lemma 6.2 with $\omega(r)=g(0)(1-s^{\prime}r)_{+}^{1/s^{\prime}}$ and $\varphi(r)=\frac{1}{s^{\prime}}(1-(g(ru)/g(0))^{s^{\prime}})$.
Therefore, we obtain that
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}g(ru)dr}{\int_{0}^{\infty}g(0)r^{\alpha-1}(1-s^{\prime}r)_{+}^{1/s^{\prime}}dr}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
for $\alpha>0$ and
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}(g(0)-g(ru))dr}{\int_{0}^{\infty}g(0)r^{\alpha-1}((1-s^{\prime}r)_{+}^{1/s^{\prime}}-1)dr}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
for $1<\alpha<0$, where we use (6.1) and (6.3). Note that
$\Big(\alpha B(\alpha,1+\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=\Big(\frac{1}{s^{\prime}}B(\alpha+1,\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=c_{n,\alpha}(s).$
Lemma 6.2 then implies that
$c_{n,\beta}(s){\rm R}_{\beta}f\subset c_{n,\alpha}(s){\rm R}_{\alpha}f,$ (6.7)
for $1<\alpha<\beta$.
Since Lemma 5.7 ensures that ${\rm R}_{\beta}f\to{\rm R}_{\infty}f$ and $c_{n,\beta}(s)\to 1$ as $\beta\rightarrow\infty$, the left-most inclusion follows directly from (6.7) by letting $\beta\rightarrow\infty$.
By Lemma 5.6 and $c_{n,\alpha}(s)\sim\frac{\alpha+1}{n+1/s}$ when $\alpha\rightarrow-1$, we have
$\lim_{\alpha\rightarrow-1^{+}}c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u)=(n+\frac{1}{s})\|f\|_{1}\rho_{\Pi^{*}f}(u),$
which yields the right-most inclusion. ∎
The following theorem demonstrates the sharpness of (6.6).
Theorem 6.9.
Let $s>0$. If $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $a=\|f\|_{\infty}>0$ and $\triangle\subset\mathbb{R}^{n}$ is a simplex containing $x_{0}$, then
${\rm R}_{\infty}f=c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f=(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$
where $1<\alpha<\beta<\infty$ and $c_{n,\alpha}(s)$ is given in Theorem 6.8.
Proof.
By Lemma 5.3, ${\rm R}_{\alpha}f$ is invariant under scaling and translation. We may assume, without loss of generality, that $x_{0}=o$ and $a=1$. Then
$\|f\|_{1}=\int_{0}^{1}|\{f\geq t\}|dt=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}|\{f^{s}\geq t\}|dt,$
where $\{f^{s}\geq t\}=(1-t)\triangle$. Hence
$\|f\|_{1}=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}(1-t)^{n}|\triangle|dt=B(\frac{1}{s},n+1)\frac{|\triangle|}{s}.$
Moreover, we obtain
$g(ru)$ $=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$
$=\int_{0}^{1}|\{f\geq t\}\cap(\{f\geq t\}+ru)|dt$
$=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}|(1-t)\triangle\cap((1-t)\triangle+ru)|dt$
$=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}(1-t)^{n}\Big|\triangle\cap(\triangle+\frac{r}{1-t}u)\Big|dt.$
By [18, Lemma 3.4], $|\triangle\cap(\triangle+ru)|^{1/n}$ is an affine function with respect to $r>0$. Therefore, there exists $\lambda>0$ such that
$\Big|\triangle\cap(\triangle+\frac{r}{1-t})u\Big|=|\triangle|\Big(1-\lambda\frac{r}{1-t}\Big)_{+}^{n},$
where $\lambda$ depends on $u$.
A straightforward calculation shows that
$g(ru)=\frac{1}{s}B(\frac{1}{s},n+1)|\triangle|(1-\lambda r)_{+}^{\frac{ns+1}{s}}.$
We set $s^{\prime}=s/(ns+1)$. For $1<\alpha<0$ and $\alpha>0$, by taking $\omega(r)=g(o)(1-s^{\prime}r)_{+}^{1/s^{\prime}}$ and $\varphi(r)=\frac{1}{s^{\prime}}(1-(g(ru)/g(o))^{s^{\prime}})$ in Lemma 6.2, we obtain
$c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u)=\frac{1}{s^{\prime}}\zeta(\alpha)=\frac{1}{s^{\prime}}\Big(\frac{\int_{0}^{\infty}r^{\alpha-1}(1-\lambda r)_{+}^{1/s^{\prime}}dr}{\int_{0}^{\infty}r^{\alpha-1}(1-s^{\prime}r)_{+}^{1/s^{\prime}}dr}\Big)^{1/\alpha}=\frac{1}{\lambda},$
where the first equality is from the calculation in the proof of Theorem 6.8.
This implies that
$c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f,\quad\quad-1<\alpha<\beta<\infty.$
Letting $\alpha\rightarrow-1$ and $\beta\rightarrow\infty$ completes the proof. ∎
To establish the equality characterization, we first need an auxiliary lemma based on the equality case of the Borell–Brascamp–Lieb inequality (2.4). Before stating this result, we collect some preliminary facts that will be used later in the argument. Throughout, ${\rm cl}K$ denotes the closure of a set $K$.
Lemma 6.10.
Let $f:\mathbb{R}\rightarrow[0,\infty)$ be a concave function on $(0,c)$ with ${\rm supp}(f)=[0,c]$, and define
$K={\rm cl}\{(x,y)\in\mathbb{R}^{2}:x\in(0,c)~\text{and}~0\leq y\leq f(x)\}.$
Let $e_{1}=(1,0)\in\mathbb{R}^{2}$. For all $r\in(0,c)$, if either
$K\cap(K-re_{1})=(1-\tfrac{r}{c})K\quad\text{or}\quad K\cap(K+re_{1})=(1-\tfrac{r}{c})K+re_{1},$
then $K$ is a triangle.
Proof.
We assume that $K\cap(K-re_{1})=(1-\frac{r}{c})K$ and prove the statement, the other case is similar. Let
$x_{0}=\inf\{x\in[0,c]:f(x)=\|f\|_{\infty}\}~\text{and}~x_{1}=\sup\{x\in[0,c]:f(x)=\|f\|_{\infty}\}.$
We first aim to show that $x_{0}=x_{1}$. If $x_{1}>x_{0}$, we choose $r>0$ to be sufficiently small such that $K\cap(K-re_{1})\neq\emptyset$ and $r<x_{1}-x_{0}$. Then the identity $K\cap(K-re_{1})=(1-\frac{r}{c})K$ implies that
$x_{1}-x_{0}-r=(1-\tfrac{r}{c})(x_{1}-x_{0}).$
It follows that either $x_{1}-x_{0}=c$ or $x_{1}-x_{0}=0$. Since the case $x_{1}-x_{0}=c$ would imply that $K$ is a rectangle, contradicting the assumption $K\cap(K-re_{1})=(1-\frac{r}{c})K$, we conclude that $x_{0}=x_{1}$.
Next we show that $f(0)=0$ if $x_{0}\neq 0$. Since $f$ is concave and hence continuous on $(0,c)$, we have $f(0^{-})=\lim_{x\rightarrow 0^{-}}f(x)=0$ and $f(0^{+})=\lim_{x\rightarrow 0^{+}}f(x)\geq f(0)\geq 0$. Thus, it suffices to show that $f(0^{+})=0$. Since $x_{0}$ is the unique maximizer, $f$ is monotone increasing on $(0,x_{0})$, and therefore $f(r)\geq f(0^{+})$ for $r>0$. Then
$\big\{(0,y):y\in[0,f(0^{+})]\big\}\subset K\cap(K-re_{1})=(1-\tfrac{r}{c})K,$
with $r$ sufficiently small. Consequently, $f(0^{+})=(1-\frac{r}{c})f(0^{+})$, which implies that $f(0)=0$. Similarly, one obtains $f(c)=0$ provided that $x_{0}\neq c$.
We now prove that $K$ is a triangle provided that $x_{0}\in(0,c)$. In this case, we have $f(0)=f(c)=0$. Since $x_{0}$ is the unique maximizer, $f$ is monotone increasing on $(0,x_{0})$ and monotone decreasing on $(x_{0},c)$. The identity $K\cap(K-re_{1})=(1-\frac{r}{c})K$ then implies that
$f\big((1-\tfrac{r}{c})x_{0}\big)=(1-\tfrac{r}{c})f(x_{0}),$
and
$f\big((1-\tfrac{r}{c})x_{0}+r\big)=(1-\tfrac{r}{c})f(x_{0}).$
This shows that $f$ is linear on $0,x_{0}]$ and affine on $x_{0},c]$, which implies that $K$ is a triangle. For $x_{0}=0$ or $x_{0}=c$, the proof is similar. ∎
Next we aim to extend Lemma 6.10 to concave functions on $\mathbb{R}^{n}$. For a non-negative, concave function $f\in L^{1}(\mathbb{R}^{n})$, we denote by
${\rm H}(f)={\rm cl}\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}~{\rm supp}(f)~\text{and}~0\leq l\leq f(x)\}$ (6.8)
the hypograph of $f$. Here ${\rm int}K$ denotes the interior of a set $K$. Let $e_{1},\ldots,e_{n}$ be an orthonormal basis of $\mathbb{R}^{n}$. To simplify the notation, we only consider the function whose support is the standard simplex defined by
$\triangle_{n}=\Big\{(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}:x_{i}\geq 0~\text{for}~i=1,\ldots,n,~\text{and}~\sum_{i=1}^{n}x_{n}\leq 1\Big\}.$ (6.9)
Recall that $a^{+}=\max\{a,0\}$. Here we also denote $a^{-}=\max\{-a,0\}$.
For $u=(u_{1},\ldots,u_{n})\in\mathbb{S}^{n-1}$, we denote by $u^{+}=(u_{1}^{+},\ldots,u_{n}^{+})$ and $u^{-}=(u_{1}^{-},\ldots,u_{n}^{-})$. Moreover, we set
$a_{0}=\sum_{i=1}^{n}u_{i}^{+},~b_{0}=\sum_{i=1}^{n}u_{i}^{-},~\text{and}~c_{0}=\max\{a_{0},b_{0}\}.$ (6.10)
One can verify geometrically that
$\triangle_{n}\cap(\triangle_{n}+ru)=(1-c_{0}r)\triangle_{n}+ru^{+},$
for any $u\in\mathbb{S}^{n-1}$.
The next result extends Lemma 6.10 to higher dimensional cases. Since we identify $\mathbb{R}^{n}$ with the hyperplane $\mathbb{R}^{n}\times\{0\}\subset\mathbb{R}^{n+1}$, we sometimes write $x\in\mathbb{R}^{n}$ also for the corresponding point in $\mathbb{R}^{n+1}$, instead of $(x,0)$.
Lemma 6.11.
Let $u\in\mathbb{S}^{n-1}$ be fixed, and $c_{0}$ be defined in (6.10). Suppose $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is concave and ${\rm supp}(f)=\triangle_{n}$, which is the standard simplex given by (6.9). For all $r\in(0,\frac{1}{c_{0}})$ and $u\in\mathbb{S}^{n-1}$, if
${\rm H}(f)\cap\big({\rm H}(f)+ru\big)=(1-c_{0}r){\rm H}(f)+ru^{+},$ (6.11)
then $f(x)=a(1-\|x-x_{0}\|_{\triangle_{n}-x_{0}})_{+}$ with $a=\|f\|_{\infty}$ and $x_{0}\in\triangle_{n}$.
Proof.
We may assume $\|f\|_{\infty}=1$ without loss of generality. We aim to show that $f$ is linear on $o,x_{0}]$ and affine on $x_{0},e_{i}]$ for all $i\in\{1,\ldots,n\}$, where $[x,y]$ denotes the line segment between $x$ and $y$. The case where $x_{0}=o$ or $x_{0}=e_{i}$ is analogous.
We first consider $f$ on $o,x_{0}]$. Let $u=-\frac{x_{0}}{|x_{0}|}$ and thus we have $u^{+}=o$. We choose the $2$-dimensional subspace $T$ spanned by $u$ and $(0,\ldots,0,1)\in\mathbb{R}^{n+1}$. Then (6.11) implies that
$K\cap(K+ru)=(1-c_{0}r)K,$
where $K={\rm H}(f)\cap T$. By Lemma 6.10, we obtain the linearity of $f$ on $o,x_{0}]$.
Next we consider $f$ on $x_{0},e_{i}]$ for any fixed $i\in\{1,\ldots,n\}$. In this case, we let $u=\frac{x_{0}-e_{i}}{|x_{0}-e_{i}|}$ and then
$u^{+}=\frac{1}{|e_{i}-x_{0}|}(x_{0}^{(1)},\ldots,x_{0}^{(i-1)},0,x_{0}^{(i+1)},\ldots,x_{0}^{(n)})~\text{and}~c_{0}=\frac{1-x_{0}^{(i)}}{|x_{0}-e_{i}|}.$
We choose the $2$-dimensional subspace $T$ spanned by $u$ and $(x_{0},1)$. Restricting (6.11) to $T$, we obtain
$\big({\rm H}(f)\cap T\big)\cap\big(({\rm H}(f)\cap T)+ru\big)=(1-c_{0}r){\rm H}(f)\cap T+ru^{+}.$ (6.12)
Set $K={\rm H}(f)\cap T\subset T$. The factor $(1-c_{0}r)$ in (6.12) corresponds to a dilation in $T$, and hence
$K\cap(K+ru)=(1-c_{0}r)\cdot K+ru.$
Here, $(1-c_{0}r)\cdot K$ denotes the dilation of $K$ in $T$ with respect to a chosen point $e_{i}$ in $T$, which serves as the origin for the scaling in the affine subspace $T$. By Lemma 6.10, $f$ is affine on $x_{0},e_{i}]$.
Therefore, we have that ${\rm H}(f)$ is a simplex in $\mathbb{R}^{n+1}$ with vertices $o,e_{1},\ldots,e_{n},(x_{0},1)$ and attains 0 on the boundary point, that is,
$f(x)=(1-\|x-x_{0}\|_{\triangle_{n}-x_{0}})_{+},$
which concludes the proof. ∎
We also need the following basic, but non-trivial result.
Lemma 6.12.
Suppose $A,B\subset\mathbb{R}^{n}$ are convex sets with ${\rm int}A\cap{\rm int}B\neq\emptyset$. Then ${\rm cl}(A\cap B)={\rm cl}A\cap{\rm cl}B$.
Proof.
Since $A\cap B\subset A$ and $A\cap B\subset B$, taking closures yields ${\rm cl}(A\cap B)\subset{\rm cl}A\cap{\rm cl}B$. Let $x\in{\rm cl}A\cap{\rm cl}B$. We choose $x_{0}\in{\rm int}A\cap{\rm int}B$ and let
$x_{\lambda}=(1-\lambda)x_{0}+\lambda x,~\text{for}~\lambda\in[0,1).$
By [42, Lemma 1.1.9], we have $x_{\lambda}\in{\rm int}A\cap{\rm int}B$. Since $x_{\lambda}\rightarrow x$ as $\lambda\rightarrow 1^{-}$, we obtain $x\in{\rm cl}({\rm int}A\cap{\rm int}B)\subset{\rm cl}(A\cap B)$. ∎
The following result associates the equality characterization of the Borell-Brascamp-Lieb inequality with (6.11) in Lemma 6.11.
Lemma 6.13.
Let $u\in\mathbb{S}^{n-1}$ be fixed, and $c_{0}$ be defined in (6.10). Suppose $f$ is a non-negative, non-zero, concave function on $\mathbb{R}^{n}$ with ${\rm supp}(f)=\triangle_{n}$. If
$\min\{f((1-c_{0}r)x+ru^{+}),f((1-c_{0}r)x+ru^{-})\}=(1-c_{0}r)f(x),$ (6.13)
for all $x\in{\rm int}\triangle_{n}$, then
${\rm H}(f)\cap\big({\rm H}(f)+ru\big)=(1-c_{0}r){\rm H}(f)+ru^{+},$
for $r\in(0,\frac{1}{c_{0}})$, where ${\rm H}(f)$ denotes the hypograph of $f$ as defined in (6.8).
Proof.
Let $c_{r}=1-c_{0}r$. For clarity of notation, we denote by
$P=\{(x,l)\in{\mathbb{R}^{n+1}}:x\in{\rm int}\triangle_{n},~0\leq l\leq f(x)\},$
and $m(x)=\min\{f((1-c_{0}r)x+ru^{+}),f((1-c_{0}r)x+ru^{-})\}$. Then ${\rm H}(f)={\rm cl}P$.
We first have
$\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}\triangle_{n},~0\leq l\leq c_{r}f(x)\}=\phi_{r}P,$ (6.14)
where $\phi_{r}={\rm diag}(1,\ldots,1,c_{r})\in{\rm GL}(n+1)$.
To consider the hypograph of $m(x)$, we first denote
$P_{1}$ $=\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}\triangle_{n},~0\leq l\leq f((c_{r}x+ru^{+})\}$ (6.15)
$~=\psi_{r}(\{(x-ru^{+},l):x\in{\rm int}\triangle_{n},~0\leq l\leq f(x)\})$
$~=\psi_{r}(P-ru^{+}),$
where $\psi_{r}={\rm diag}(c_{r}^{-1},\ldots,c_{r}^{-1},1)$. Similarly, we denote
$P_{2}=\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}\triangle_{n},~0\leq l\leq f((c_{r}x+ru^{-})\}=\psi_{r}(P-ru^{-}).$ (6.16)
Note that $\psi_{r}=c_{r}^{-1}\phi_{r}$ and $(A-x)\cap(A-y)=A\cap(A+(x-y))-x$ holds for any set AA and $x,y$. Combining this with (6.14), (6.15), (6.16) and the assumption (6.13), we obtain
$P\cap(P+ru)=c_{r}P+ru^{+}.$ (6.17)
Recall that $r\in(0,\frac{1}{c_{0}})$, which yields that ${\rm int}(P+ru)\cap{\rm int}P\neq\emptyset$. Then Lemma 6.12 implies that
${\rm cl}\big(P\cap(P+ru)\big)={\rm cl}P\cap({\rm cl}P+ru).$
Since ${\rm cl}P={\rm H}(f)$, we conclude the proof by taking closures on both sides of (6.17). ∎
Combining Lemma 6.13 with Lemma 6.11, we obtain the following crucial corollary.
Corollary 6.14.
For $u\in\mathbb{S}^{n-1}$, let $c_{0}$ be defined in (6.10). Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a non-zero, concave function with ${\rm supp}(f)=\triangle_{n}$. For $r\in(0,\frac{1}{c_{0}})$, if
$\min\{f((1-c_{0}r)x),f((1-c_{0}r)x+ru)\}=(1-c_{0}r)f(x),$
for all $x\in{\rm int}\triangle_{n}$ and $u\in\mathbb{S}^{n-1}$, then $f(x)=a(1-\|x-x_{0}\|_{\triangle_{n}-x_{0}})_{+}$ with $a=\|f\|_{\infty}$ and $x_{0}\in\triangle_{n}$.
We are now in a position to prove the following key result for the equality characterization in Theorem 3.
Lemma 6.15.
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, $s$-concave function with compact support. For $u\in\mathbb{S}^{n-1}$, let
$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx.$
If there is a constant $\lambda>0$ such that $g(ru)=\|f\|_{1}(1-\lambda r)_{+}^{\frac{ns+1}{s}}$ for $r>0$, then $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
Proof.
We assume $\|f\|_{\infty}=1$ without loss of generality. Since g(ru)=g(ru)g(ru)=g(-ru), we set
$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x-ru)\}dx,$
and $N_{f}(x,r)=\min\{f(x),f(x-ru)\}$ to simplify notation. Let $r\geq 0$ and $\lambda\in[0,1]$ be fixed, and define the functions $F,G,H:\mathbb{R}^{n}\to\mathbb{R}$ by
$F(x)=N_{f}(x,0)=f(x),~~G(x)=N_{f}(x,r),~~H(x)=N_{f}(x,\lambda r),$
which then satisfy
$H((1-\lambda)x+\lambda y)\geq\Big((1-\lambda)F(x)^{s}+\lambda G(y)^{s}\Big)^{1/s}=\mathcal{M}_{\lambda}^{s}(F(x),G(y))$
for all $x,y\in\mathbb{R}^{n}$ .
Since $r\mapsto g(ru)^{\frac{s}{ns+1}}$ is an affine function with respect to $r$, then
$\int_{\mathbb{R}^{n}}H=\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}F,\int_{\mathbb{R}^{n}}G\Big).$
Hence, equality holds in the Borell-Brascamp-Lieb inequality (2.4) for $F,G,H$.
By Theorem 2.4, there exists $c_{r}=\Big(\frac{|{\rm supp}(G)|}{|{\rm supp}(f)|}\Big)^{1/n}$ and $y_{r}\in\mathbb{R}^{n}$ such that
${\rm supp}(G)=c_{r}\,{\rm supp}(f)+y_{r},$ (6.18)
while the definition of $G$ gives
${\rm supp}(G)={\rm supp}(f)\cap({\rm supp}(f)+ru).$ (6.19)
Since $u\in\mathbb{S}^{n-1}$ is arbitrarily fixed, (6.18) and (6.19) imply that ${\rm supp}(f)\cap({\rm supp}(f)+ru)$ is homothetic to ${\rm supp}(f)$ for all $u\in\mathbb{S}^{n-1}$. By [18, Lemma 3.4], ${\rm supp}(f)$ is a simplex denoted by $\triangle$.
By Lemma 5.3, ${\rm R}_{\alpha}f$ is invariant under translations and ${\rm R}_{\alpha}(f\circ\phi^{-1})=\phi{\rm R}_{\alpha}f$ for all $\phi\in{\rm GL}(n)$. It suffices to consider the case where $\triangle$ is the standard simplex, that is,
$\triangle=\triangle_{n}=\Big\{x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}:x_{i}\geq 0,~~\sum_{i=1}^{n}x_{i}\leq 1\Big\}.$
By (6.18) and (6.19), we have
$c_{r}=1-c_{0}r,~~\text{and}~~y_{r}=ru^{+},$
where $c_{0}$ is defined by (6.10).
Moreover, by Theorem 2.4, there exists a $(\lambda_{0},s)$-concave function $\Phi:{\rm supp}(F)\rightarrow[0,\infty)$, with $\lambda_{0}=\frac{\lambda c_{0}}{1-\lambda+\lambda c_{0}}$, satisfying
$\left\{\begin{aligned} &F(x)=\Phi(x)&&,\\ &G(c_{r}x+y_{r})=c_{r}^{\frac{1}{s}}\Phi(x)&&,\\ &H((1-\lambda+\lambda c_{0})x+\lambda y_{r})=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{r}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&,\end{aligned}\right$.
for almost all $x\in\triangle_{n}$. By the definition of $F$, we have $\Phi(x)=F(x)=f(x)$, and
$G((1-c_{0}r)x+y_{r})^{s}=\min\{f((1-c_{0}r)x+ru^{-})^{s},f((1-c_{0}r)x+ru^{+})^{s}\}=(1-c_{0}r)f(x)^{s}.$ (6.20)
Since $f^{s}$ is concave and hence continuous on ${\rm int}\triangle_{n}$, the equality (6.20) holds for all $x\in{\rm int}\triangle_{n}$. By applying Corollary 6.14, we complete the proof. ∎
As a consequence, Lemma 6.15 derives the following theorem, which provides the full equality characterization in Theorem 3.
Theorem 6.16.
Let $s>0$ and $c_{n,\alpha}(s)$ be as defined in Theorem 3. Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. For $1<\alpha<\beta<\infty$, if
$c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f,$
then $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
Proof.
Let $g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$ for a fixed $u\in\mathbb{S}^{n-1}$ and $r\in\mathbb{R}$. Setting $\omega(r)=g(o)(1-s^{\prime}r)_{+}^{1/s^{\prime}}$ and $\varphi(r)=\frac{1}{s^{\prime}}(1-(g(ru)/g(o))^{s^{\prime}})$ in Lemma 6.2, we obtain
$\zeta(\alpha)=s^{\prime}c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u),$
where $s^{\prime}=s/(ns+1)$. Then, the equality $c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f$ implies that $\zeta(\alpha)=\zeta(\beta)$. Since $\omega$ is strictly decreasing on its support, Lemma 6.3 yields $\lambda=\zeta(\alpha)^{-1}$ such that
$\varphi(r)=\frac{1}{s^{\prime}}\Big(1-\Big(\frac{g(ru)}{g(o)}\Big)^{s^{\prime}}\Big)=\lambda t.$
This shows that $g(ru)^{s^{\prime}}$ is affine in $r$, and the conclusion follows from Lemma 6.15. ∎
We conclude this section by showing the Rogers-Shephard inequality and Zhang inequality for $s$-concave functions, which are special cases of Theorem 3.
Theorem 6.17 (Rogers-Shephard inequality for $s$-concave functions).
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. Then
$|{\rm D}\,{\rm supp}(f)|\leq\frac{\Pi_{k=1}^{n}(k+\frac{1}{s})}{n\cdot n!\|f\|_{1}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy.$
Equality holds if and only if $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
Theorem 6.18 (Zhang’s inequality for $s$-concave functions).
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. Then
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy\leq\frac{n\cdot n!(n+\frac{1}{s})^{n}}{\Pi_{k=1}^{n}(k+\frac{1}{s})}\|f\|_{1}^{n+1}|\Pi^{*}f|.$
Equality holds if and only if $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
Acknowledgements: This research was funded in whole or in part by the Austrian Science Fund (FWF) doi/10.55776/37030. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.